### v1

# Optimal or Greedy Decision Trees? Revisiting their Objectives, Tuning, and Performance

Abstract
Decision trees are traditionally trained using greedy heuristics that locally optimize an impurity or information metric. Recently there has been a surge of interest in optimal decision tree (ODT) methods that globally optimize accuracy directly. We identify two relatively unexplored aspects of ODTs: the objective function used in training trees and tuning techniques. Additionally, the value of optimal methods is not well understood yet, as the literature provides conflicting results, with some demonstrating superior out-of-sample performance of ODTs over greedy approaches, while others show the exact opposite. In this paper, we address these three questions: what objective to optimize in ODTs; how to tune ODTs; and how do optimal and greedy methods compare? Our experimental evaluation examines 13 objective functions, including four novel objectives resulting from our analysis, seven tuning methods, and six claims from the literature on optimal and greedy methods on 165 real and synthetic data sets. Through our analysis, both conceptually and experimentally, we discover new non-concave objectives, highlight the importance of proper tuning, support and refute several claims from the literature, and provide clear recommendations for researchers and practitioners on the usage of greedy and optimal methods, and code for future comparisons.
Keywords: optimal decision trees, CART, objectives, complexity tuning, classification
1 Introduction
Decision trees (DTs) are among the most-used (interpretable) machine learning (ML) models. Despite their simplicity, they can learn complex non-linear relationships in data. Their human comprehensibility answers the need for interpretable models in high-stake domains (Rudin, 2019; Arrieta et al., 2020), provided the trees are small. Optimal decision trees (ODTs) specifically, which provably optimize an objective for a given size limit, provide small but accurate models on many tabular datasets and thus combine high performance with interpretability (Piltaver et al., 2016; Loh, 2014; Carrizosa et al., 2021).
Because training optimal decision trees with respect to a size limit is NP-Hard (Hyafil and Rivest, 1976), most early decision tree learning methods were greedy top-down induction heuristics, such as CART (Breiman et al., 1984) and C4.5 (Quinlan, 1993), which locally optimize some impurity or information gain metric for each branching node.
Consequently, greedy decision tree learning has been extensively studied given its long history; important examples being splitting criteria (Mingers, 1989b; Buntine and Niblett, 1992; Shih, 1999; Raileanu and Stoffel, 2004; Wang and Xia, 2017) and pruning techniques to avoid overfitting (Mingers, 1989a; Patil et al., 2010; Esposito et al., 1997).
In contrast, optimal decision tree research is a much younger field, with the last decade seeing major advancements. The main topic of research has been improving scalability by reducing runtimes and supporting larger datasets. Researchers have employed a variety of techniques such as (mixed) integer programming (MIP) (Bertsimas and Dunn, 2017; Verwer and Zhang, 2017), constraint programming (Verhaeghe et al., 2020), satisfiability (Hu et al., 2020; Narodytska et al., 2018) and dynamic programming with bounds (Aglin et al., 2020a; Lin et al., 2020; Demirović et al., 2022). Whereas ten years ago optimal methods were limited to small data sets, due to algorithmic advancements and an increase in computation power, we can now use a recent dynamic programming approach (van der Linden et al., 2023) to analyze data sets with up to hundreds of thousands of samples.
Given that the main technical innovations for optimal methods are relatively recent, unlike in the field of greedy heuristics, principled ways of using optimal decision trees for out-of-sample performance have been comparatively under-explored. Whereas the field of greedy decision trees shows a large variety of splitting criteria, pruning methods, and tuning approaches, optimal decisions are almost exclusively trained by maximizing accuracy, possibly additionally penalizing the number of nodes (the sparse objective). Tuning, if done at all, is performed in different ad hoc manners, e.g., tuning the number of nodes or depth of the tree. Practices differ from paper to paper, which hinders direct comparisons.
Moreover, early comparisons between optimal and greedy approaches were limited in scope and contained claims and hypotheses that we can now refute (Section 5). Murthy and Salzberg (1995) lacked a scalable ODT method and therefore confined their analysis on synthetic data. Bertsimas and Dunn (2017) trained ODTs using MIP, but lack of scalability constrained most of their analysis to data sets of only 100 instances or trees with a maximum depth of two. For larger problems, their approach did not converge to optimality; therefore, the support for several of their claims remained uncertain.
Though these and other studies (Lin et al., 2020; Demirović et al., 2022) report an average improvement of the out-of-sample performance versus greedy heuristics, others have criticized ODTs for overfitting (Dietterich, 1995), observed worse results for ODTs compared to greedy heuristics (Zharmagambetov et al., 2021; Marton et al., 2024), and questioned the adjective ‘optimal’ (Sullivan et al., 2024). These contradictory findings illustrate the need for a more thorough understanding of the concept of optimal decision trees.
This motivated us to conduct a thorough review of existing decision tree methods, both greedy and optimal, focusing on objective functions used during training and different tuning approaches. Based on our findings, we motivate new objective functions and tuning approaches that are specific to optimal methods that globally optimize the objective. To support our study, we conducted the largest evaluation to this date concerning optimal and greedy decision tree methods, taking into account 13 different objective functions, seven tuning approaches, 165 real-world and synthetic data sets (small and large), and trees that go beyond small tree-depth limits. This provides us with a wealth of data to analyze and insights into the strengths and weaknesses of greedy and optimal approaches.
From our new insights obtained on training ODTs, we also discuss the implications for decision tree learning in general. To keep the scope of this study manageable, we chose to limit this study to axis-aligned binary classification trees with hard splits, which are arguably the most common type of decision trees. In more detail, we contribute the following:
  • In Section 3, we analyze and compare nine existing greedy decision tree accuracy objectives for optimal decision trees. We observe that the strict concavity of many of these objectives, as required by greedy top-down inducting approaches (Kearns and Mansour, 1996), yields counter-intuitive results when trained to optimality. Since ODTs do not require concave objectives, we contribute four new non-concave objectives. Our experiments show that these new non-concave objectives outperform the default accuracy objective for smaller data sets by more effectively penalizing leaf nodes with nearly uniform class distributions. For larger data sets, ODTs can directly optimize the in-sample accuracy.
  • In Section 4, we compare seven complexity tuning approaches for ODTs, four of which were proposed before, and three new tuning approaches. Our experiments highlight the importance of tuning optimal decision tree methods, and that (surprisingly) the differences between the commonly used tuning methods are small.
  • In Section 5, we review previous comparisons between greedy and optimal approaches, formulate best practices for future comparisons, and provide data and code to support proper benchmarking. [^1] The code will be made public on paper publication. We apply these practices in evaluating six claims from the literature on the performance of greedy and optimal trees. Specifically, we evaluate the following claims:
    1. Claim 1:
      Optimal methods under the same depth constraint (up to depth four) find trees with 1-2% higher out-of-sample accuracy than greedy methods (Bertsimas and Dunn, 2017; Verwer and Zhang, 2017; Demirović et al., 2022).
    2. Claim 2:
      Optimal methods have a better accuracy-interpretability trade-off than greedy methods (Lin et al., 2020).
    3. Claim 3:
      The difference between optimal and greedy approaches diminishes with more data (Murthy and Salzberg, 1995; Costa and Pedreira, 2023).
    4. Claim 4:
      The accuracy of greedy trees remains stable when the data size increases linearly with concept complexity (Murthy and Salzberg, 1995).
    5. Claim 5:
      Optimal trees are more likely to overfit than greedy trees (Dietterich, 1995).
    6. Claim 6:
      The expected depth of greedy trees remains (in practice) close to the optimal tree depth, even for complex (large) true tree size (Goodman and Smyth, 1988; Murthy and Salzberg, 1995).
    Our results support Claims 126, and refute Claims 34 and 5.
The remainder of the paper is organized as follows. The next section provides a general overview of the field of optimal decision trees. Sections 3, 4, and 5 are mostly self-contained sections, each dedicated to a single major research question as outlined above, each with its corresponding related work, technical details, experiments, and conclusions. Section 6 draws an overarching conclusion.
2 Related Work
This section introduces previous literature on decision tree learning with a focus on optimal methods. For longer reviews, we refer to surveys by Safavian and Landgrebe (1991); Kotsiantis (2013) and Costa and Pedreira (2023).
Decision tree learning started several decades ago with AID (Morgan and Sonquist, 1963), a recursive approach to regression analysis, later adapted for classification in CHAID (Kass, 1980). Since then, two of the most popular decision tree learning algorithms have been CART (Breiman et al., 1984) and ID3 (Quinlan, 1986) with its successors C4.5 (Quinlan, 1993). Each of these uses top-down induction (TDI) to greedily partition the data by finding a split that is locally optimal according to an information or impurity criterion. Overfitting is prevented either by early stopping rules such as a minimum information gain or by post-pruning the tree. We discuss splitting criteria in Section 3.1.
Besides TDI heuristics, other heuristics include stochastic gradient descent (Norouzi et al., 2015), coordinate descent (Carreira-Perpinán and Tavallali, 2018; Dunn, 2018; Bertsimas and Dunn, 2019), evolutionary algorithms (Barros et al., 2011; Guidotti et al., 2024), swarm optimization (Panhalkar and Doye, 2022), and look-ahead (Kiossou et al., 2024). These metaheuristics typically obtain better trees by considering a larger search space than TDI heuristics but do not guarantee to find the globally optimal tree.
Because computing trees optimally is NP-Hard (Hyafil and Rivest, 1976; Cox et al., 1989; Murphy and McCraw, 1991), historically most approaches have been heuristics. Although a few early optimal dynamic programming (Schumacher and Sevcik, 1976; Payne and Meisel, 1977; Miyakawa, 1985; Cox et al., 1989; Nijssen and Fromont, 2007, 2010) and an extreme point tabu search (Bennett and Blue, 1996) approaches have been proposed, only recently, with increasing computation power and new algorithmic improvements, interest in computing optimal trees resurged.
This resurge started with mixed-integer programming (MIP) formulations for ODTs (Bertsimas and Dunn, 2017; Verwer and Zhang, 2017) with several consecutive improvements (Verwer and Zhang, 2019; Aghaei et al., 2021; Liu et al., 2024; Günlük et al., 2021; Zhu et al., 2020; Hua et al., 2022; Alès et al., 2024). The advantages of these MIP methods are that they can find splits with arbitrary thresholds on the continuous features and can easily be adapted by adding linear constraints or changing the objective, including objectives that operate on the whole tree instead of summing the objectives of independent leaf nodes. The disadvantage is poor scalability because of a weak linear relaxation and the inability of the MIP solver to recognize the independence between subtrees.
Around the same time (maximum) satisfiability (SAT) formulations were proposed (Narodytska et al., 2018; Hu et al., 2020; Janota and Morgado, 2020; Avellaneda, 2020; Shati et al., 2023; Alòs et al., 2023). These SAT models focus on finding perfect trees of minimum size and MaxSAT is used to maximize the training accuracy for a fixed size limit. Similarly, Verhaeghe et al. (2020) find perfect trees of minimum size using constraint programming. They improve performance by exploiting the subtree independence, by caching and reusing solutions to subproblems, and by pruning the search through bounds.
Similar techniques are exploited in dynamic-programming (DP) based approaches, which a recent survey (Costa and Pedreira, 2023) indicates as the most promising approach in terms of scalability. DP exploits the independent subtree structure and reuses partial solutions to repeated subproblems. The addition of bound-based pruning (Aglin et al., 2020a, b), improved lower bounds (Hu et al., 2019; Lin et al., 2020; Demirović et al., 2022; Chaouki et al., 2024), and a faster subprocedure for trees of depth two (Demirović et al., 2022) have greatly improved the scalability of the basic DP approach. The advantages of DP are the good scalability for realistic use cases, specifically with respect to the number of instances. The disadvantages are the need for binarization and an exponential runtime with respect to the number of features and maximum tree size.
Recent developments for ODTs are incorporating continuous features (Mazumder et al., 2022; Shati et al., 2023); exploring the Rashomon set of all close to optimal models (Xin et al., 2022; Semenova et al., 2023); quantifying and reducing explanation redundancy (Izza et al., 2022; Audemard et al., 2022); improving anytime performance (Kiossou et al., 2022; Demirović et al., 2023); improving memory usage (Aglin et al., 2022); and applying ODTs to other objectives such as regression (Zhang et al., 2023; van den Bos et al., 2024), quantile regression (Lemaire et al., 2024), fairness constraints (Aghaei et al., 2019; Jo et al., 2022; van der Linden et al., 2022), robustness (Vos and Verwer, 2022; Justin et al., 2022), survival analysis (Zhang et al., 2024; Huisman et al., 2024), prescriptive policy generation (Bertsimas et al., 2019; Jo et al., 2021; van der Linden et al., 2023), and learning MDP policies (Vos and Verwer, 2023).
In summary, the recent literature shows a surge in methods for and applications of ODTs. Advances in scalability make it now possible to do a more in-depth analysis of how ODTs should be trained and how they compare to the traditional greedy approaches.
3 The Optimization Objective for Optimal Decision Trees
Decision tree learning objectives typically optimize two parts: some accuracy objective (i.e., accuracy or one of its proxies, such as information gain) and a tree-complexity objective (e.g., number of nodes). Some objectives only include the first, while others, such as the minimum description length (see below), combine both into one. In this section, we focus on the first: the accuracy objective. In Section 4, we discuss the tree-complexity objective.
Section 3.1 lists existing greedy splitting criteria and how they relate to ODT objectives. In our analysis of these objectives in Section 3.2, we observe that the concavity traditionally required by greedy heuristics (Kearns and Mansour, 1996) is not helpful when training ODTs. Therefore, we introduce four novel non-concave objectives in Section 3.3. We then empirically compare all these objectives in Section 3.4 and discuss our findings in Section 3.5. Our main finding is that optimal and greedy decision tree learners benefit from optimizing different objectives. Greedy learners perform best with concave objectives while optimal learners achieve the best performance when optimizing non-concave objectives.
3.1 From Greedy Criteria to Objectives
Greedy TDI criteria have been extensively analyzed and compared. We extend this comparison to optimal decision trees. In the following, we first consider how existing greedy criteria relate to objective functions for ODTs.
Greedy top-down induction (TDI) methods typically consider two types of criteria: a local splitting criterion that decides how to split a node and a pruning criterion that decides which nodes to keep during pruning. Although the goal of decision tree learning is to maximize accuracy, TDI methods rarely directly optimize accuracy but instead optimize a proxy. One of the first classification tree methods, CHAID (Kass, 1980), maximizes the reduction in the $\chi^{2}$ statistic; CART (Breiman et al., 1984) optimizes the Gini impurity; and ID3 (Quinlan, 1986) and C4.5 (Quinlan, 1993) optimize information gain (or entropy). Some greedy methods use the same criterion for splitting and pruning; others employ another criterion for pruning such as the minimum description length (Mehta et al., 1995) or the pessimistic error rate in C4.5 (Quinlan, 1993). We explain each in more detail below.
ODTs, on the other hand, do not have a splitting or pruning criterion but optimize a global objective (also called a loss function), typically expressed as a sum over the errors of the leaf nodes plus some term that penalizes the size of the tree. Almost all ODT methods directly optimize the accuracy (Nijssen and Fromont, 2007; Bertsimas and Dunn, 2017; Verwer and Zhang, 2017, 2019; Hu et al., 2019; Aglin et al., 2020a; Hu et al., 2020; Aghaei et al., 2021). To improve the performance on unbalanced data sets, some have considered other objectives such as balanced accuracy (Lin et al., 2020) or F1-score (Demirović and Stuckey, 2021). Nijssen and Fromont (2010) have implemented the pessimistic error objective from C4.5 and a Bayesian objective function, but do not discuss the effect on the out-of-sample performance when these objectives are used.
We now present a list of splitting and pruning criteria used for TDI heuristics and for each provide an equivalent formulation as an ODT leaf node objective. We rewrite these functions as a function $f$ with as input the leaf node size $n$ and the number of misclassifications $e$. Let $L$ be a list of leaf nodes of size $n$ and with $e$ misclassifications, then the minimization objective of the whole tree becomes
$\sum_{(n,e)\in L}f(n,e)\,.$
For brevity, we consider only the binary classification case. For all functions, terms that divide by zero (e.g., $e/n$ when $n=0$) or take the log of zero (e.g., $\log_{2}(e/n)$ when $e=0$) are evaluated as zero. We leave out some splitting criteria, such as the twoing rule in CART, if they are not additive or cannot easily be expressed as an objective function for the leaf.
Accuracy:
The most direct way to optimize out-of-sample accuracy is to optimize the in-sample accuracy by minimizing the misclassification score. The leaf node objective in terms of the leaf node size $n$ and the number of misclassifications $e$ is
$f(n,e)=e\,.$
A possible disadvantage of this objective is its ‘flatness’: it cannot distinguish between many different solutions. For example, $f(4,2)+f(6,0)=f(5,1)+f(5,1)$.
Gini impurity:
Commonly used in greedy top-down heuristics is the Gini impurity. Gini impurities are from 0 to 1. Weighted Gini impurity scores are obtained by multiplying the Gini impurity by the number of instances in that leaf node. Let $p_{0}=\frac{e}{n}$ denote the probability of the first class and $p_{1}=\frac{n-e}{n}$ the probability of the second class. Then the objective value is
$f(n,e)=n(1-p_{0}^{2}-p_{1}^{2})\,.$
Square root Gini:
Kearns and Mansour (1996) propose to use the square root of the Gini impurity to improve performance on unbalanced data sets:
$f(n,e)=n\sqrt{1-p_{0}^{2}-p_{1}^{2}}\,.$
Entropy:
The second commonly used objective in top-down heuristics is entropy (or information gain). Like the Gini impurity, weighted scores are obtained by multiplying by the size of the leaf node:
$f(n,e)=-\frac{n}{2}(p_{0}\log_{2}p_{0}+p_{1}\log_{2}p_{1})\,.$
Gini impurity and entropy can be expressed as a parameterization of Tsallis entropy (Tsallis, 1988; Wang and Xia, 2017).
Minimum error:
Niblett (1987) estimates the expected error for nodes by assuming that every class has equal probability. With $n$ samples in a node and $|\mathcal{K}|$ the number of labels, and $n_{c}$ the count of the majority label, the expected error rate is
$\frac{n-n_{c}+|\mathcal{K}|-1}{n+|\mathcal{K}|}\,.$
Since we consider binary classification, $|\mathcal{K}|=2$, and $n_{c}=n-e$. Therefore:
$f(n,e)=\frac{n(e+1)}{n+2}$
This is equivalent to Laplace smoothing with a smoothing parameter set to one (add-one smoothing) (Flach, 2012). According to Mingers (1989a), the equal-probability assumption of this approach becomes problematic for a large number of classes.
Pessimistic error:
Quinlan (1987) proposed a pessimistic error rate by computing a bound on the expected error rate, which in effect raises the error rate at every leaf by 0.5:
$f(n,e)=e+\frac{1}{2}\,.$
This method is similar to a complexity cost per node of 0.5. Since we cover complexity costs in the next section, we do not consider the pessimistic error in this section.
Binomial pessimistic error (Binom.):
The commonly used C4.5 method (Quinlan, 1993) uses an advanced form of pessimistic error by considering a leaf with $n$ training instances and $e$ misclassifications as a ‘sample’ from a binomial distribution with an unknown misclassifying probability. Since this probability cannot be computed directly, the upper confidence bound of the posterior distribution of this probability, based on a confidence level $\alpha$, is used as the error rate of the leaf node. The confidence interval width $z_{\alpha}$ is the $z$-value from the normal distribution for confidence level $\alpha$. Let $e^{\prime}=e+\frac{1}{2}$ be the pessimistic error. Then the binomial pessimistic error can be expressed as:
$f(n,e)=\begin{cases}n\cdot\left(1-\operatorname{exp}^{\ln{(\alpha)}/n}\right)&\text{if\leavevmode\nobreak\ }e=0\\ e&\text{if\leavevmode\nobreak\ }e=n\\ \frac{e^{\prime}+\frac{z_{\alpha}^{2}}{2}+\sqrt{z_{\alpha}^{2}\left(e^{\prime}\left(1-\frac{e^{\prime}}{n}\right)+\frac{z_{\alpha}^{2}}{4}\right)}}{n+z_{\alpha}^{2}}\cdot n&\text{otherwise.}\end{cases}$
Minimum description length (Quinlan):
The minimum description length (MDL) approach states that the best model can be described with the least amount of bits of information because the description length of a model can directly be linked to the posterior probability of a model (Rissanen, 1978; Li and Vitányi, 2008). In practice, the encoding typically consists of two parts: first the encoding of the model and then the encoding of the data that deviates from the model.
Quinlan and Rivest (1989) observe that the cost of encoding a binary string of length $n$ with $e$ ones and $n-e$ zeros can be computed by first encoding the size $n$ of the string, and then the positions of the $e$ ones, with $b$ representing an upper bound on the number of ones that can occur:
$L(n,e,b)=\ln{(b+1)}+\ln{\binom{n}{e}}\,.$ (1)
Then, for every leaf node with $n$ instances and $e$ misclassifications, encode a bit string that specifies the misclassifications with cost $L(n,e,\left\lfloor{n/2}\right\rfloor)$ for binary classification:
$f(n,e)=L(n,e,\lfloor(n+1)/2\rfloor)\,.$
The encoding of the branching feature and the leaf node label are part of the tree complexity cost which we cover in the next section.
Minimum description length (Mehta):
Mehta et al. (1995) observe a sub-optimal coding length for Eq. (1) when $e$ is close to zero, and therefore propose to use stochastic complexity (Krichevsky and Trofimov, 1981; Rissanen, 1997). For binary classification, their formula can be rewritten to
$f(n,e)=e\ln\frac{n}{e}+(n-e)\ln\frac{n}{n-e}+\frac{1}{2}\ln\frac{n}{2}+\ln\pi\,.$
Bayesian:
Decision trees are also commonly trained using a Bayesian approach (Chipman et al., 1998; Denison et al., 1998). These approaches find the maximum likelihood tree given some priors. We present the objective function used in the recent work by Sullivan et al. (2024). For the binary case, they assume that each leaf node can be represented by a Bernoulli distribution with parameter $\theta\in[0,1]$. They assume $\theta\in\operatorname{Beta}(\rho_{0},\rho_{1})$, the Beta distribution with parameters $\rho_{0},\rho_{1}\in\mathbb{R}^{+}$. The values $\rho_{0}$ and $\rho_{1}$ are hyperparameters, but they fix these values to $\rho_{0}=\rho_{1}=2.5$. The error can then be expressed in terms of the Beta function $B$ as follows:
$f(n,e)=\frac{B(e+\rho_{0},n-e+\rho_{1})}{B(\rho_{0},\rho_{1})}\,.$
Sullivan et al. (2024) also add a cost based on the priors, which can be considered a complexity cost, and therefore we leave it out in the discussion in this section.
3.2 Analysis
Figure 1: (Left) Three splitting heuristics compared. The horizontal axis shows the binary class distribution expressed as the probability of the first class, and the vertical axis shows the corresponding splitting criterion value (lower is better). (Right) Geometric interpretation of the weighted mean error of two children when $p$, $p_{1}$, and $p_{2}$ represent the class distributions of the parent and the two children respectively. The length of the arrow indicates the improvement in the splitting criterion value. Adapted from Flach (2012).
The reason why TDI methods do not optimize accuracy directly is that an accuracy splitting criterion is often unable to find an improving split in unbalanced data. When splitting a node, TDI methods evaluate all possible splits and choose the split that minimizes the splitting criterion value for the resulting class distributions among the new nodes of each possible split. Fig. 1 visually explains why using accuracy as a splitting criterion is worse at distinguishing improving splits than Gini impurity or entropy. The left side shows the function values for accuracy, Gini impurity, and entropy. The right side shows how a locally optimal split can be found geometrically. When splitting a node with a probability of the first class of $p$ into two new nodes with probabilities $p_{1}$ and $p_{2}$, the new weighted splitting criterion value can be found by drawing a straight line from the criterion value at point $p_{1}$ to $p_{2}$. The intersection of the straight line at $p$ is the sum of the weighted criterion value of the two nodes. For Gini impurity and entropy, this value is always lower than the criterion value of the parent node, because both functions are strictly concave. Accuracy, however, is not strictly concave, and when $p<0.5,p_{1}<0.5$, and $p_{2}<0.5$ (or equivalently all are greater than $0.5$), the weighted sum of the criterion value of the child nodes is the same as that of the parent node. Moreover, for any values $p_{1}<0.5$ and $p_{2}<0.5$ the weighted sum of the criterion values is the same, and therefore no distinction can be made between these splits. Thus TDI heuristics require strictly concave splitting criteria (Kearns and Mansour, 1996) and therefore do not optimize accuracy directly.
Figure 2: Objective values for different objective functions for a single leaf node. Left, the leaf node size is fixed at $n=100$. Right, the misclassifications are fixed at $e=20$. Surprisingly, the value of the strictly concave objectives increases for a fixed error and increasing leaf node size.
To further understand the differences between all the objectives listed above, we plot them and analyze the differences. The left of Fig. 2 shows the values of the objectives presented above when the leaf node size is fixed but the number of misclassifications increases. The accuracy is a straight line since every misclassification is counted equally. Both the pessimistic binomial score and the minimum error follow the accuracy tightly, with only a small additional cost for higher misclassifications. All other objectives follow roughly the same pattern: the first misclassifications in a node are penalized most and the additional penalty for extra misclassifications decreases.
Similarly, on the right of Fig. 2, when the number of misclassifications in a node is fixed but the leaf node size is changed, the accuracy is a straight line. The pessimistic binomial score and the minimum error again follow accuracy closely. Interestingly, for all other objectives, the objective value increases when the node size increases. Since lower values are preferred, this means these objectives penalize larger leaf nodes more than smaller leaf nodes with the same misclassifications.
Intuitively, it seems odd that the objective increases for larger nodes with a fixed error. Table 1 shows some examples of relative objective values that are unexpected. For example, according to entropy, it is better to have two nodes of size 4, with two misclassifications in the first node and zero in the second, than one node of size 8 with one misclassification. Entropy strongly values a pure node, even if this means a higher misclassification rate. Other objectives, such as MDL, value two nodes of size four and two with two misclassifications in the first and none in the second, more than one node of size six with also two misclassifications. Again, a small pure node is valued, even though the node of size four with two misclassifications has a high probability of being misclassified.
Table 1: Pure nodes are overvalued, resulting in splits with pure nodes (e.g., $[n,e]=[4,0]$) and nodes that are labeled randomly ($[4,2]$), rather than keeping one node with the same misclassifications ($[8,2]$) or even less in case of entropy ($[8,1]$).
Objective Expected (lower is better) Observed
Gini impurity $f(8,2)<f(4,2)+f(4,0)$ $f(8,2)=3.000$, $f(4,2)+f(4,0)=2.000$
Entropy $f(8,1)<f(4,2)+f(4,0)$ $f(8,1)=2.174$, $f(4,2)+f(4,0)=2.000$
MDL (Quinlan) $f(6,2)<f(4,2)+f(2,0)$ $f(6,2)=4.708$, $f(4,2)+f(2,0)=4.377$
MDL (Mehta) $f(6,2)<f(4,2)+f(2,0)$ $f(6,2)=5.513$, $f(4,2)+f(2,0)=5.409$
Bayes $f(6,2)<f(4,2)+f(2,0)$ $f(6,2)=4.379$, $f(4,2)+f(2,0)=4.321$
3.3 Novel Non-Concave Objectives
The odd behavior of the greedy criteria in Fig. 2 and Table 1 is a result of their strict concavity. (Strict) concavity is not a requirement for ODTs because ODTs do not consider splitting criteria and can search beyond a non-improving split. Therefore we here introduce four non-concave objectives that have not previously been used in decision tree learning.
Error Gini:
Instead of weighing the Gini impurity by the leaf node size, we propose Error Gini which multiplies the Gini impurity by the number of errors in a leaf node. This ensures that the objective value does not increase when the misclassifications are fixed but the size of the leaf node increases. Again, let $p_{0}=\frac{e}{n}$ and $p_{1}=\frac{n-e}{n}$, then the objective value is
$f(n,e)=e(1-p_{0}^{2}-p_{1}^{2})\,.$
Error entropy:
Similarly, we provide Error entropy which gives the entropy multiplied by the number of errors, instead of weighted by the size of the leaf node:
$f(n,e)=-e(p_{0}\log_{2}p_{0}+p_{1}\log_{2}p_{1})\,.$
M-loss:
Noel et al. (2023) propose two new objective functions for optimizing neural networks, both of which are non-concave. The first new objective function they propose is the M-loss, here rewritten in terms of $n$ and $e$:
$f(n,e)=n\left(\frac{1}{1-\frac{e}{n}}-1\right)\,.$
L-loss:
The second objective function that they propose is called the L-loss. Rewritten in terms of $n$ and $e$ this becomes
$f(n,e)=n\left(\frac{1}{\sqrt{1-\left(\frac{e}{n}\right)^{2}}}-1\right)\,.$
Figure 3: The new objectives show opposite behavior to the strictly concave objectives. Left, the leaf node size is fixed at $n=100$. Right, the misclassifications are fixed at $e=20$.
Fig. 3 shows the values these new functions take. For easy comparison, accuracy and Gini impurity are also included in the plot. In contrast to the strictly concave functions, the left side of Fig. 3 shows how the first misclassifications in a leaf node are penalized less, whereas nodes with a (close to) balanced class distribution are heavily penalized. The right side shows that increasing the leaf node size while keeping the number of misclassifications constant, decreases the penalization. Therefore, these objectives obtain the desired property to penalize nodes with a close-to-equal class distribution more strongly.
3.4 Experiments
In our experiments we aim to answer the following questions:
  1. 1.
    What is the difference between the objectives on out-of-sample accuracy when trained to optimality on the training data?
  2. 2.
    Which objectives are well suited for small data sets?
  3. 3.
    What difference can be observed between the objectives when trained to optimality on the training data or when greedily optimized using TDI heuristics?
3.4.1 Experiment Setup
We empirically compare the objectives on a large benchmark set from OpenML (Vanschoren et al., 2013; Feurer et al., 2021). [^2] The code will be made public on paper publication. For the sake of scalability, we selected all binary classification data sets with 50 or fewer features, of which ten or fewer numeric features and no large text features, with no missing values, with at most 100,000 instances, and at least 40 instances. We take the most recent version of the data set and omit duplicates or data sets that only differ in the random seed, resulting in 165 data sets. We split each data set into five folds, creating five train and test pairs each consisting of four and one fold respectively. We list all data sets used in Appendix A.
We implemented all the ODT objectives in the ODT method STreeD (van der Linden et al., 2023) because of its scalability and flexibility in supporting new objectives. [^3] https://github.com/algtudelft/pystreed STreeD is a DP approach that requires binary features. Therefore we binarize the numeric training data with thresholds on the ten quantiles, and the categorical data with one-hot encoding (with at most ten categories). The test data is binarized in the same way. We experimented with other binarization approaches but noted no significant impact. See Section 5.2 for the impact of binarization on CART.
We train ODTs up to a depth of three and four while tuning both the tree depth and the number of branching nodes using five-fold cross-validation. A depth-three tree has at most eight leaf nodes and seven branching nodes and a depth-four tree has at most 16 leaf nodes and 15 branching nodes.
We test the objectives on TDI heuristics in our own implementation of CART. We use cost-complexity tuning with accuracy as the pruning objective. For comparison with ODT, we train CART using a depth limit of four on the same binarized data sets. In Section 5.2, we evaluate the impact of these choices.
[…]
Figure 4: Comparing ODT objectives for (a) $\operatorname{max-depth}=3$ and (b) $\operatorname{max-depth}=4$. Orange (blue) indicates (non-)concave. The average rank, accuracy, and number of leaf nodes over all data sets and folds are shown.
Figure 5: Nemenyi critical distance rank test for ODTs of maximum depth four. The average rank per objective is plotted and objectives with a rank difference smaller than the critical distance (CD) at p-value $0.05$ are grouped by a black bar.
3.4.2 Optimal Decision Tree Results
Figs. 4 and 4 show the average rank, test accuracy, and the number of leaf nodes per objective for trees of depth three and four. On average, the best ranking objectives with the best test accuracy are accuracy and its slight variations Minimum Error and the Binomial Pessimistic error, and our novel non-concave functions M-loss, error entropy, error Gini, and L-loss. Maximizing the training accuracy is the second and highest-ranked objective for maximum depths three and four respectively. However, the highest average test accuracy for both depth limits is obtained by minimizing the new objective M-loss. For a maximum depth of four, the pessimistic binomial objective achieves a similar average test accuracy with a lower number of leaf nodes than the accuracy objective.
However, the results are close. Fig. 5 shows the results of a Nemenyi critical distance rank test to test the significance (Demšar, 2006). This test computes the critical distance (CD) between the average ranks of two methods to be statistically significant. Fig. 5 shows that all the non-concave objectives and Gini are not significantly different for depth four. The non-concave objectives are significantly better than the concave objectives (except Gini).
3.4.3 Small Data Sets
We noted that the differences among objectives were larger for smaller data sets. These data sets are particularly hard because overfitting is likely with such limited information. Therefore Fig. 6 shows the results on the 70 data sets with less than 200 instances.
Figure 6: Comparing ODT objectives for $\operatorname{max-depth}=4$ on 70 small data sets ($n<200$).
The new non-concave objectives perform significantly better than accuracy, with the test accuracy of M-loss on average exceeding that of the accuracy objective by $0.7\%$. Therefore, with sufficient data, ODTs obtain the best performance by maximizing training accuracy but for smaller data sets, other objectives, such as the M-loss can yield better results.
3.4.4 Greedy Heuristics Results
Figure 7: Comparing greedy objectives for $\operatorname{max-depth}=4$. The strictly concave objectives outperform the non-concave objectives.
For comparison, we also train greedy trees with the objectives listed above. Fig. 7 shows the resulting out-of-sample performance. In comparison with the ODTs, the ranking of the objectives is almost reversed, which is in line with our analysis in Section 3.2. A Nemenyi critical distance rank test shows that all strictly concave objectives are significantly better than all non-concave objectives for TDI heuristics. The smaller tree sizes of the non-concave objectives show that the greedy heuristic gets stuck early with these objectives and finds no improving splits. Optimizing Gini impurity yields 1.2% higher out-of-sample accuracy than directly optimizing the in-sample accuracy. These results confirm that the traditional Gini and entropy are among the top choices.
3.5 Discussion
Previous work has extensively compared greedy splitting criteria, which we extend to optimal decision trees. Our results show that optimizing accuracy directly is a good choice for ODTs, specifically when the training data is large because, with sufficient training data, the training accuracy approximates the test accuracy. This shows that the strict concavity of objectives such as Gini impurity and entropy, is not an inherently necessary or desirable property, but a limitation imposed by the greedy TDI approach. Optimal and greedy training procedures respond differently to the objectives and therefore best practices of one approach do not necessarily translate to the other.
We obtained better results for small data sets by using one of our new non-concave objectives, such as M-loss or error entropy. These objectives strongly penalize leaf nodes with a close to equal class distribution and are therefore more robust when there is little data available. This does not apply to greedy heuristics which show an almost opposite ranking of objectives in comparison to ODTs. This opens the question of how the performance of non-concave objectives for small data sets could be exploited in greedy heuristics.
4 Tuning the Tree Complexity of Optimal Decision Trees
Most decision tree learning approaches make a trade-off between training accuracy and model complexity to prevent overfitting. For ODTs, several complexity tuning methods have been used but without an in-depth empirical comparison. Therefore this section analyses the effect of complexity tuning methods for optimal decision trees, starting with an overview of complexity tuning methods. We find that optimal decision trees perform significantly better with tuning than without, but that existing tuning techniques perform similarly.
4.1 Tuning Approaches
The following complexity tuning approaches are currently used in ODT methods:
Depth:
A common metric for tree complexity is its depth. DL8.5 (Aglin et al., 2020a, b), for example, tunes the depth of the tree. A possible disadvantage is that the number of nodes increases exponentially with respect to the maximum depth, thus providing only a coarse control of the tree size. In our experiments below, we tune the depth parameter $d\in\{0,...,\operatorname{max-depth}\}$, with a $\operatorname{max-depth}$ of either four or five.
Size:
The number of nodes can be tuned directly as a hard constraint. Since binary trees always have one more leaf node than branching nodes, the total number of nodes can also be counted by the number of branching or leaf nodes alone. This approach is taken by Demirović et al. (2022) and van der Linden et al. (2023) who directly tune both the maximum depth and the number of branching nodes. In our experiments below, we tune the depth and the number of nodes such that the depth $d\in\{0,...,\operatorname{max-depth}\}$ and the number of branching nodes $n\in\{d,...,2^{d}-1\}$.
Complexity cost:
The most common approach in greedy trees is to penalize the cost of adding a node by a factor $\lambda$. This approach is used in most MIP methods (Bertsimas and Dunn, 2017) and also in the optimal sparse approaches by Lin et al. (2020) and Chaouki et al. (2024). Complexity-cost tuning minimizes $\lambda(|L|-1)+\sum_{(n,e)\in L}f(n,e)$, with $L$ the set of leaves and $\lambda$ the complexity-cost parameter.
As similarly done in (Lin et al., 2020), we tune the complexity-cost parameter $\lambda\in\{0.0001$, $0.0002$, $0.0003$, $0.0004$, $0.0005$, $0.0006$, $0.0007$, $0.0008$, $0.0009$, $0.001$, $0.002,0.003$, $0.004$, $0.005$, $0.006$, $0.007$, $0.008$, $0.009$, $0.01$, $0.02$, $0.03$, $0.04,0.05$, $0.06$, $0.07$, $0.08$, $0.09$, $0.1$, 0.2}0.2\}0.2 } in our experiments below.
Minimum Support:
The minimum support or minimum leaf node size is a hard constraint on the minimum number of instances that should end up in a leaf node. Closely related is the minimum split size: the minimum number of instances in a node to allow for a split. Minimum support is more common in the data mining literature, e.g., Nijssen and Fromont (2007). It is also introduced as a soft constraint (Vilas Boas et al., 2021).
In our experiments below, we tune the minimum leaf node size $m$ with possible values for $m$ expressed as a percentage $m^{\prime}$ of the data set size $n$ with $m^{\prime}\in\{0.001$, $0.002$, $0.003$, $0.004$, $0.005$, $0.0075$, $0.01$, $0.015$, $0.02$, $0.03$, $0.04$, $0.05$, $0.06$, $0.07$, $0.08$, $0.09$, $0.1$, $0.15$, $0.2$, $0.25$, $0.3$, $0.35$, $0.4$, $0.45$, 0.5}0.5\}0.5 }. Rounding $m^{\prime}n$ to an integer may yield non-unique numbers, so we select the smallest ten unique integer possible values for $m$.
In our experiments, we leave out selecting a tree structure from a fixed set of possible structures (Günlük et al., 2021), because this approach was chosen specifically to deal with the scalability limitations of MIP.
4.2 New Tuning Approaches
On top of the existing ODT tuning approaches, we evaluate three new approaches:
Question length:
[…]
G-Test threshold:
The G-Test threshold approach measures the difference between the two branches of a split and only allows splits that result in splits with a statistically significant difference. The G-Test is a likelihood ratio test that works analogously to the $\chi^{2}$ test used in CHAID (Kass, 1980) and ID3 (Quinlan, 1986). In our experiments below, we tune the significance level $\alpha\in\{0.000001$, $0.00001$, $0.0001$, $0.001$, $0.005$, $0.01$, $0.05$, $0.1$, $0.2$, $0.5$, 1.0}1.0\}1.0 }.
Smoothing:
Because of the good performance of the minimum error objective in Section 3.1, we generalize this approach by tuning the Laplace smoothing parameter. The Laplace smoothing approach (Flach, 2012) assumes in a leaf node that for each class, $x$ extra instances exist. With $|\mathcal{K}|$ the number of classes, the accuracy objective becomes
$f(n,e)=\frac{n(e+x)}{n+|\mathcal{K}|x}\,.$
We tune the smoothing parameter $x\in\{0$, $1$, $2$, $3$, $4$, $5$, $7$, $10$, $15$, 20}20\}20 }.
4.3 Experiments
[…]
Figure 8: Complexity tuning results for optimal decision trees of (a) $\operatorname{max-depth}=4$ and (b) $\operatorname{max-depth}=5$ for five runs on (a) 164 data sets and (b) 131 data sets.
Figure 9: Nemenyi critical distance rank test for ODTs of maximum depth four. The average rank per tuning method is plotted and methods with a rank difference smaller than the critical distance (CD) at p-value $0.05$ are grouped by a black bar.
In our experiments we aim to answer the following question: how do the ODT tuning methods compare in out-of-sample accuracy?
We tune each method using five-fold cross-validation on the 165 OpenML sets data described in Section 3.4.1.444The code will be made public on paper publication. Regardless of the tuning method, we impose a maximum depth constraint of four or five. For runtime comparison, we ran all experiments on an Intel Xeon E5-6248R 3.0GHz with 100GB RAM using one thread. All tuning methods were implemented in STreeD (van der Linden et al., 2023).
Figs. 8 and 8 show the performance of the complexity tuning method on the OpenML data sets. We exclude data sets from the analysis if any method exceeded the one-hour time out: for depth four, one data set out of 165 is removed, and for depth five, 34 data sets. The results show that tuning the size performs best, although all methods are close.
Surprisingly, Fig. 9 shows that tuning only the depth is not statistically significantly worse even though it tests much fewer configurations than other approaches. On average it does yield larger trees since ODT methods will use as many nodes as allowed to maximize the objective on the training data. Tuning the minimum support also yields larger trees. In our experiments, tuning the question length works well for depth four but worse for depth five, but it is unclear why. Tuning through smoothing is also a new promising approach. Tuning the G-Test threshold and no tuning are significantly worse than tuning the size.
Inspecting the results shows that the differences among tuning methods are largest for medium-sized data sets (several hundred to ten thousand samples). For example, on the 77 data sets with 250 or more and 10,000 or fewer samples, a Wilcoxon signed rank test shows tuning the size is statistically significantly better than tuning the depth (p-value $\approx 0.03$). From this, we hypothesize that for small data sets the validation sets are too small to yield good estimates of the test scores, whereas for large data sets, the risk of overfitting and thus the importance of tuning decreases.
Fig. 10 shows the runtime for a five-fold cross-validation over the parameter options described above for trees of maximum depth four. Note that these runtimes could easily be changed by considering more, less, or other possible parameter configurations, although this may also impact the accuracy performance. The left plot shows what percentage of computations for the five runs on the 165 data sets finished within the given runtime. The middle and right plots show the runtime for data sets for an increasing number of instances and binary features. Each marker represents the average runtime for one data set.
Apart from no tuning, which yields significantly worse accuracy results, tuning the depth or the size of the tree is the most efficient approach because of how STreeD, the ODT method used in the experiments, works. Since STreeD is a dynamic programming approach, it can reuse previously cached results when searching for increasingly large trees.
Figure 10: Runtime distributions for each tuning methods for $\operatorname{max-depth}=4$. The scatter plots show one marker for every data set averaged over five runs. Tuning the depth or the size of the tree is the most efficient tuning approach.
In conclusion, complexity tuning of ODTs is necessary. All previously used tuning approaches obtain similar accuracy results. Tuning the depth or the size of the tree (when using cached solutions) is more efficient than other approaches. Two new tuning approaches are promising: tuning the question length and smoothing. Future work could further explore combining multiple approaches, e.g., tuning the size and the smoothing level.
Based on our experiments in Sections 3 and 4, we recommend to train optimal decision trees as follows:
{mdframed}
Recommendations 1 (Training Optimal Decision Trees).
  1. 1.
    Optimize the same loss at train and test time for optimal decision trees.
    Optimizing the accuracy on average yields the best out-of-sample accuracy.
  2. 2.
    For small data sets, maximize a non-concave objective such as M-loss.
    For data sets with less than a few hundred samples, optimizing non-concave objectives such as M-loss performs better than maximizing accuracy as it has an added regularization effect.
  3. 3.
    Tune the complexity of optimal decision trees.
    Training ODTs with hyperparameter tuning is significantly better than training without hyperparameter tuning.
  4. 4.
    Tune the size, complexity cost, or smoothing parameter; or, in the case of large data sets, the depth.
    Tuning the size, complexity cost, smoothing parameter, or depth, on average, yields similar out-of-sample results. However, tuning the depth yields larger trees and is worse for medium-sized data sets (a few hundred to ten thousand samples). Using DP approaches, tuning the size is more runtime efficient. For large data sets, tuning is less important, and tuning the depth is more runtime efficient.
5 Optimal Compared with Greedy
To understand the differences between greedy and optimal approaches for learning decision trees, we collected claims from the literature and evaluated these claims with extensive experiments. Below, we list the claims discussed in this section and summarize their verdicts:
Claim 1:
Optimal methods under the same depth constraint (up to depth four) find trees with 1-2% higher out-of-sample accuracy than greedy methods (Bertsimas and Dunn, 2017; Verwer and Zhang, 2017; Demirović et al., 2022).
Supported:
We evaluate the accuracy of depth three and four trees on 164 data sets and find an average improvement of 1.1% and 1.3% of optimal over greedy approaches.
Claim 2:
Optimal methods have a better accuracy-interpretability trade-off than greedy methods (Lin et al., 2020).
Supported:
We evaluate the accuracy of trees with 1 to 16 nodes on 165 data sets and find that the size-weighted accuracy of optimal methods is, on average, higher than that of greedy methods. The size-weighted accuracy is a new metric we propose in Section 5.3 to measure the accuracy-interpretability trade-off.
Claim 3:
The difference between optimal and greedy approaches diminishes with more data (Murthy and Salzberg, 1995; Costa and Pedreira, 2023).
Refuted:
Experiments on synthetic and real data show that the performance of size-unconstrained greedy trees keeps improving with more samples, while also growing much larger trees than ODTs. In contrast, size-constrained greedy trees do not improve after some point and stay worse than optimal.
Claim 4:
The accuracy of greedy trees remains stable when the data size increases linearly with concept complexity (Murthy and Salzberg, 1995).
Refuted:
On synthetic data generated from a random decision tree, the performance of optimal decision trees remains stable when the random tree’s size is increased, and the performance of greedy trees diminishes.
Claim 5:
Optimal trees are more likely to overfit than greedy trees (Dietterich, 1995).
Refuted:
With hyperparameter tuning, we do not find significant performance differences between optimal and greedy methods with small numbers of samples (up to 200) nor more sensitivity to noise. When trained with the same size constraint, optimal methods dominate with more samples.
Claim 6:
The expected depth of greedy trees remains (in practice) close to the optimal tree depth, even for complex (large) true tree size (Goodman and Smyth, 1988; Murthy and Salzberg, 1995).
Supported:
Our experiments on synthetic data with ground-truth trees of depths 1 to 5 show that the expected depth of all methods remains similar to the true depth.
In this section, we first review previous greedy-optimal comparisons, which we use to design a set of best practices for future comparisons. The rest of the section details the experiments we performed to evaluate each of the claims from existing literature. In Section 5.5 we discuss the scalability of ODTs.
5.1 Previous Comparisons
This section reviews previous comparisons between optimal and greedy decision tree learning methods, each different in its experiment design and sometimes in its conclusion. We restrict our review to binary axis-aligned trees. First, we summarize the comparisons from papers that proposed ODT methods. Second, we discuss other papers that compare greedy heuristics and optimal methods. Finally, we discuss the best and poor comparison practices as input for a fair comparison method presented after. We highlight the encountered claims that were mentioned above.
5.1.1 Comparisons in ODT Papers
Papers that propose new ODT methods typically aim to train a decision tree with a given size constraint that achieves the best out-of-sample performance. Nijssen and Fromont (2007, 2010) compare their optimal DL8 algorithm with J48, an implementation of C4.5. When trained on the same discretized data, without a depth limit, but with the same minimum support constraint, DL8 is significantly better for 9 out of 20 data sets and worse for one, while yielding trees that are 1.5 times larger than J48. However, when J48 is trained without the minimum support constraint and with the non-discretized data, J48 outperforms DL8 on out-of-sample accuracy for most data sets.
Bertsimas and Dunn (2017) develop the optimal MIP method OCT and compare it with CART on real and synthetic data. When CART is constrained to the same depth limit as OCT (up to depth four), they conclude that OCT, on average, has a statistically significant 1-2% better out-of-sample accuracy (Claim 1). The largest difference is observed at depth two. They hypothesize that the smaller difference for depths three and four is the result of OCT not reaching optimality within their time limit. When CART is run with a depth limit of ten, it is negligibly better than OCT at depth four.
The main restriction of their analysis is the scalability of OCT. Because of this, they restrict synthetic data analysis to data sets with only 100 instances and two continuous features. They also experiment with data sets up to 1600 instances, but only on ground truth trees of depth two. When training OCT on the synthetic data, they set the maximum depth to the true depth, which prevents overfitting.
Verwer and Zhang (2019) compare the optimal MIP methods BinOCT, DTIP (Verwer and Zhang, 2017), and OCT with depth-constrained CART on data sets with a few thousand instances. They report results without hyperparameter tuning and observe that the ODTs are significantly better for depths two and three and slightly better for depth four (Claim 1).
Lin et al. (2020) propose GOSDT, an ODT method with a sparsity coefficient. They conclude that GOSDT obtains a better accuracy-interpretability trade-off than other methods, including CART (Claim 2). This is based on an experiment on six small data sets with a coarse binarization applied to both GOSDT and CART. CART is tuned using the maximum depth parameter (from one to six), instead of tuning the complexity-cost as is normally done. They tune GOSDT using complexity-cost tuning without a depth limit.
Demirović et al. (2022) compare their optimal MurTree algorithm and CART on binarized data sets with up to 43500 instances and 1163 binary features. They run MurTree for different depths (from one to four) and number of nodes, and CART for different depths (from one to four), and report the best test accuracy for each method. They too report an average out-of-sample improvement of 1-2% over CART (Claim 1).
The same trend appears in other papers that propose new ODT methods. The aim is to show ODTs’ superior performance under a fixed depth limit. An exception is (Alès et al., 2024), who compare with unconstrained greedy approaches. Results are often shown for fixed hyperparameters (Hu et al., 2020; Mazumder et al., 2022; Liu et al., 2024). Scalability limits the analysis to small data sets (Hu et al., 2020; Günlük et al., 2021; Alès et al., 2024) or larger data sets are run only with a maximum depth of two (Hua et al., 2022).
5.1.2 Other Comparisons
Papers that do not propose new ODT methods typically have another aim: the best out-of-sample accuracy without imposing depth constraints on the tree.
Murthy and Salzberg (1995) compare the greedy approach to known true (optimal) trees on a synthetic benchmark. They observe that the greedy tree is approximately one standard deviation larger than the true tree size while the expected depth (question length) is very close to optimal, which was also observed by Goodman and Smyth (1988); Murphy and McCraw (1991) (Claim 6). However, the maximum depth of greedy trees is on average approximately two times higher than the true depth. When they increase the data set size linearly with the true tree complexity, they observe almost no drop in accuracy for the greedy approach (Claim 4). From this result, Costa and Pedreira (2023) hypothesize that the gap between optimal and greedy approaches diminishes for more data (Claim 3).
Dietterich (1995) concludes from the empirical results by Quinlan and Cameron-Jones (1995) that optimal methods are more prone to overfitting (Claim 5). Exhaustively searching through all possible models may yield smaller models, but is also more prone to finding small patterns that do not represent the ground truth. Therefore Dietterich concludes that it is better to train greedy methods with a model complexity penalty.
Zharmagambetov et al. (2021) compare greedy methods with their local search method TAO (Carreira-Perpinán and Tavallali, 2018) and the optimal methods OCT and GOSDT. They conclude that most methods perform similarly, except TAO, which outperforms the other methods. In many cases, CART outperforms OCT and GOSDT. For CART and TAO, they train greedy trees up to depth 30. For OCT, they use the results reported in (Bertsimas and Dunn, 2017), which go up to depth four. They train GOSDT with a high complexity cost, yielding trees that are on average no larger than 3.4 leaf nodes for any of the data sets.
Sullivan et al. (2024) propose MAPTree, a search algorithm that finds the maximum a posteriori tree. They compare with DL8.5, GOSDT, and CART and conclude that MAPTree outperforms these approaches, which leads them to question the ‘optimality’ of ODTs. They observe that DL8.5 is prone to overfitting, while GOSDT is prone to underfitting, and both are sensitive to hyperparameter tuning, whereas MAPTree is not. However, these results are from averaging the performance per hyperparameter setting over all data sets, rather than tuning the hyperparameters for each individual run. Additionally, they evaluate MAPTree without a depth limit, while other methods (including CART) are run with a depth limit.
Marton et al. (2024) learn axis-aligned trees with gradient descent and compare the results a.o. to CART and DL8.5. GradTree outperforms the other methods for binary classification, while CART performs the best for multi-class classification. The ODT approach DL8.5 performs the second worst in both cases: only the evolutionary approach is worse. They tune each method using random search, except for DL8.5 which they fix to a maximum depth of four. The other methods were typically trained up to depths 7-10.
5.1.3 Discussion
Year Author(s) Method     Greedy with same size limit     Greedy without size limit     Small and large data sets     Beyond small trees     Optimal tuned (correctly)     Greedy tuned (correctly)
Papers that propose ODT methods
2007 Nijssen and Fromont DL8
2017 Bertsimas and Dunn OCT
2019 Verwer and Zhang BinOCT
2020 Lin et al. GOSDT
Hu et al. MaxSAT
2021 Günlük et al. ODT
2022 Demirović et al. MurTree
Hua et al. RS-OCT
Mazumder et al. Quant-BnB
2024 Liu et al. BNP-OCT
Alès et al. CTT
Other papers that compare ODTs with greedy DTs
1995 Murthy and Salzberg -
2021 Zharmagambetov et al. -
2024 Sullivan et al. MAPTree
Marton et al. GradTree
Table 2: Simplified overview of the comparisons between greedy and optimal methods in the literature. Ideally, a comparison checks all columns. 1) Compare methods under the same size constraint; 2) compare (greedy) methods without a size constraint; 3) compare on small and large data sets (> 10.000 instances); 4) compare optimal methods beyond depth three; 5) tune optimal methods (correctly); and 6) tune greedy methods (correctly).
From the comparisons listed above, the following general observations can be made about how greedy and optimal methods are compared:
  • Both greedy and optimal methods are often not correctly tuned, or not tuned at all.
  • The papers that propose ODTs typically compare ODTs with greedy methods under the same size constraint, while other papers compare ODTs with a restricting size limit and other greedy methods without (or with a very permissive) a size limit.
  • Comparisons are often limited to a small number of small data sets.
  • Comparisons are sometimes limited to small trees.
We briefly discuss each of these in more detail.
When comparing with CART, the papers above show several modifications to how CART is trained and tuned. We assess the impact of each of these modifications on the accuracy in Section 5.2 below.
Several papers evaluated ODTs without tuning the size. As shown in Section 4.3, this significantly worsens the ODT performance. Others evaluate ODTs with a substantial restriction on the tree size, resulting in shallow underfitting trees.
Almost all the ODT papers compare with CART under the same tree-size constraint and draw a positive conclusion on the ODTs’ performance. The other papers typically compare ODTs under a size constraint with unconstrained non-optimal approaches and sometimes draw negative conclusions about the ODTs’ performance. Both comparisons are useful for different purposes. The central claim from the ODT papers is that ODTs have a better accuracy-interpretability trade-off. Therefore, we further discuss measuring this trade-off in Section 5.3 below.
Several comparisons are limited to small data sets and small trees (e.g., less than 10000 instances or trees of maximum depth three). This is typically because of scalability limitations. The improved scalability of optimal methods allows us to analyze larger data sets and trees.
Table 2 summarizes our observations about previous comparisons. We recommend that future comparisons between ODTs and greedy approaches should check all the columns in this table. We summarize our recommendations as follows:
{mdframed}
Recommendations 2 (Greedy-Optimal Comparisons).
  1. 1.
    Compare both with and without constraining the sizes of the decision trees.
    Optimal decision trees optimize performance for a particular size; therefore, greedy methods should be equally constrained in size to compare fairly.
  2. 2.
    Compare performance on both small and large data sets.
    While experiments on small data are more efficient to run, their results often do not carry over to larger data sets.
  3. 3.
    Evaluate both small and large trees.
    Previous comparisons often compare trees of depth two. This optimization problem is too simplistic, and the results do not always carry to a larger depth.
  4. 4.
    Tune both greedy and optimal methods and ensure an equal comparison.
    Comparisons between greedy and optimal trees at a fixed depth are unequal since the methods respond differently to size constraints. If one wants to compare trees up to depth four, for example, both approaches should be tuned up to a maximum depth of four, not trained with a fixed depth of four.
5.2 Training CART
Figure 11: Results for CART (Gini) when trained with(out) binarization, with(out) a depth limit, and when using depth or cost-complexity tuning. Using a depth limit (as indicated in blue) significantly impacts performance. Tuning the depth has a more negative impact for large maximum depths. Binarization with 10 quantile thresholds has no significant impact on the accuracy but does impact the tree size.
In the experiments reviewed above, ODTs are often compared to a modified version of CART, for example, to allow for a direct comparison under similar circumstances. We test the impact of these modifications to assess the validity of these previous comparisons and inform future comparisons. We assess the following typical modifications: 1) tuning the depth instead of the complexity cost; 2) binarizing the feature data; or 3) running CART while imposing an additional depth constraint.
We compare CART’s performance with these modifications against unmodified CART on the 165 OpenML data sets used before. We approximate the unconstrained CART with a maximum depth of 20. We set the constrained depth limit to four, because of its common use in ODT comparisons. As before, the binarized data has up to ten binary features per continuous feature by using thresholds on ten quantiles or one-hot encoding of categorical features with a maximum of ten categories.
We apply cost-complexity pruning as done in RPart (Therneau et al., 2023): we train a fully expanded tree and obtain the cost-complexity path with all possible cost-complexity values from that tree and use the geometric mean to get the midpoints of those values. We use five-fold cross-validation to find the best cost-complexity parameter among the midpoints and retrain a tree on the full training data with this parameter.
Fig. 11 shows CART’s performance under these modifications. The largest differences are between the depth-constrained and the unlimited depth variant. Binarization has only a small impact on the performance (this does not necessarily generalize for more coarse binarizations). When a strict depth limit is imposed, tuning the depth instead of the complexity cost has a small impact but for the unlimited depth case, this significantly hurts CART’s performance. Fig. 11 also shows significantly different tree sizes for CART’s modifications. Both binarization and depth tuning result in larger trees. From these results we can conclude the following best practices:
  1. 1.
    Training CART with a depth limit should be clearly stated.
  2. 2.
    Tuning the depth of CART instead of the complexity cost should be avoided.
  3. 3.
    Training on binarized data does not significantly harm the performance, provided the binarization is not too coarse, but should be explicitly noted.
5.3 Accuracy-Interpretability Trade-Off
Figure 12: Typical accuracy-interpretability trade-off for greedy and optimal. Optimal has a slight advantage for small trees but both methods converge for large size limits.
ODT papers typically compare ODTs with greedy heuristics under a similar size constraint. This is partly motivated by the definition of ODTs because ODTs are defined as trees that maximize training performance under a given size limit. No theoretical claim is made about its out-of-sample performance or the performance without a size constraint. In fact, finding an optimal tree without a size constraint is trivial: it is obtained by splitting in any way until no further split can be made. It is precisely the size constraint that makes it NP-Hard (Hyafil and Rivest, 1976).
It is also motivated by the claim that ODTs have a better accuracy-interpretability trade-off. Since one of the oft-cited motivations for decision trees is their comprehensibility and large trees with hundreds of nodes can hardly be called human-comprehensible (Piltaver et al., 2016), evaluating an algorithm’s ability to obtain small performant trees is important.
Typically, the tree size and test accuracy are plotted against each other, as done in Fig. 12, to assess this trade-off. Such plots show the relative performance for different size limits and also the saturation point: when adding more nodes gives no improvement in accuracy. Optimal methods typically reach this saturation point earlier and greedy methods eventually catch up by obtaining the same accuracy but with larger trees.
Such comparisons require one figure per data set. To enable easier comparisons across a large number of data sets, we propose a new metric: the size-weighted accuracy (SWA). The purpose of this metric is to express the accuracy-interpretability trade-off with one number that represents the ‘surface’ under the accuracy-interpretability Pareto front seen in Fig. 12. Since we are mostly interested in the performance of small trees, we set the weight of the accuracies obtained for a tree with $i$ leaf nodes to $1/i$. We define the size-weighted accuracy ($\operatorname{SWA}_{n}$) as the weighted average of the obtained trees of maximum size $n$:
$\operatorname{SWA}_{n}=\frac{1}{\sum_{i=1}^{n}\frac{1}{i}}\sum_{i=1}^{n}\frac{\operatorname{acc}_{i}}{i}\,.$ (2)
Since not every algorithm can directly set the number of leaf nodes, we propose to run the algorithm with different complexity parameters (e.g., the complexity-cost parameter $\lambda$) and record the resulting number of leaf nodes and test accuracy for each run. If multiple runs yield the same number of leaf nodes, then average these test accuracies. For missing tree sizes, linearly interpolate test accuracies using results from the nearest smaller and larger tree sizes. If the largest tree size obtained is less than $n$, assign the test accuracy of this largest tree to all larger sizes up to $n$. If larger trees result in smaller test accuracy, replace it with the larger value for smaller trees, since we measure the Pareto front.
For example, for the synthetic data in Fig. 12, we computed trees up to 20 leaf nodes. CART obtains a $\operatorname{SWA}_{20}$ of $84.0\%$, whereas the optimal approach obtains a $\operatorname{SWA}_{20}$ of $84.7\%$.
5.4 Experiments on Literature Claims
To evaluate the claims made in previous papers, we compare optimal and greedy methods on synthetic and real data sets. Our experiment setup follows the best practices introduced in Recommendations 2.555The code will be made public on paper publication.
5.4.1 Experiment Setup
We follow the synthetic data setup from Murthy and Salzberg (1995); Bertsimas and Dunn (2017) and Dunn (2018). We generate $n$ random training instances with $p$ numeric features, uniformly distributed over $[0,1]$. For a given percentage $f$ of the training data, we add feature noise by adding noise uniformly drawn from $[-0.25,0.25]^{p}$. We binarize the numeric features by threshold predicates on 10 quantiles per numeric feature. We generate a random binary tree on this binarized data of a maximum depth $d$ with at most $2^{d}$ leaf nodes. We choose random splits on the data such that each leaf node contains at least 5 instances. The binary labels of each leaf node are assigned alternately, such that no split leads to two leaf nodes with the same label. After this, we add class noise to a given percentage $c$ of the data by flipping its label. By default, we set the true tree depth to three, the number of instances $n=1000$, the number of numeric features $p=3$, the feature noise $f=0\%$, and the class noise $c=0\%$. For each training set, we create a corresponding test set without noise of $1000$ instances per leaf node in the generated tree.
[…]
We also evaluate the methods on the same OpenML data sets as in Section 3.4 and add two more datasets: covertype and Higgs. These additional datasets were chosen for their large number of samples, allowing us to investigate performance under various sample sizes.
We evaluate both CART and ODT on the binarized data to eliminate this difference between the two methods and focus only on the difference between greedy and optimal search. Comparing on binarized data is commonly done (Lin et al., 2020; Demirović et al., 2022). We show in Section 5.2 that for the data sets in our benchmark, the impact of this binarization on CART is only small. We train ODTs using STreeD by maximizing accuracy and tuning the size of the tree, according to our analysis in Section 3 and 4 (Recommendations 1). We train CART trees with and without a depth limit, using cost-complexity tuning, following the best practices mentioned in Section 5.2.
In our synthetic experiments, apart from the familiar test accuracy and number of leaf nodes, we also measure the following:
True Discovery Rate (TDR):
The TDR is the percentage of splits in the ground truth tree that are recovered in the trained tree (higher is better).
False Discovery Rate (FDR):
The FDR is the percentage of the splits in the trained tree that are not part of the ground truth tree (lower is better).
Question Length:
The Question length (or expected depth) is the average number of branching nodes an instance visits when evaluated (lower is better).
Figure 13: Optimal (blue) versus CART (red) out-of-sample accuracy results on five runs for 164 data sets. CART without a depth limit performs similarly to optimal with a maximum depth of four. CART with the same depth limit performs significantly worse.
Figure 14: Nemenyi critical distance rank test for optimal versus CART. The average rank per method is plotted and methods with a rank difference smaller than the critical distance (CD) at p-value $0.05$ are grouped by a black bar.
5.4.2 Out-of-Sample Accuracy
Claim 1.
Optimal methods under the same depth constraint (up to depth four) find trees with 1-2% higher out-of-sample accuracy than greedy methods (Bertsimas and Dunn, 2017; Verwer and Zhang, 2017; Demirović et al., 2022).
Figure 15: The difference between the optimal and CART out-of-sample accuracy for both small (left) and larger (right) data sets. The dashed line indicates the average difference.
To evaluate Claim 1, we evaluate both the ODT approach and CART on the OpenML data sets with a depth limit of three and four. We also compare with CART without a depth limit. Since ODT with a maximum depth of four hits the time-out of one hour for one data set, we report the results for the other 164 data sets.
Fig. 13 and 14 show that ODT and greedy (without a depth limit) perform similarly on average, but the greedy approach yields much larger trees. When compared under the same depth constraint as stated in Claim 1, optimal outperforms greedy with an average improvement of 1.1% and 1.3% for depths three and four respectively. These differences are statistically significant. Since optimal algorithms optimize the complete decision tree, instead of greedily improving the tree, they can achieve better scores.
Fig. 15 shows the distribution of the differences between optimal and greedy when trained with maximum depth four. For small data sets ($n\leq 250$), the average advantage of the optimal approach is $0.7\%\pm 8.3$. Because data is sparse, both methods are sensitive to overfitting, resulting in some large accuracy differences, but on average, the optimal approach is better. For larger data sets ($n>250$), the average improvement of optimal over CART is $1.9\%\pm 4.7$. The difference is larger and the variance is lower. These results confirm Claim 1.
5.4.3 Accuracy-interpretability trade-off
Optimal ($\operatorname{max-depth}=4$)
Data set $n$ $p$ Binom. Accuracy CART
Space_ga 3107 6 68.7 68.7 67.5
Phoneme 3172 5 69.8 69.7 68.4
Pollen 3848 5 50.6 50.6 50.5
Analcatdata_supreme 4052 7 91.0 90.9 90.8
Abalone 4177 8 69.4 69.4 68.7
Wilt 4839 5 96.4 96.4 96.2
Banana 5300 2 69.8 69.8 69.5
Delta_ailerons 7129 5 81.8 81.7 81.7
Bank8fm 8192 8 80.1 80.0 79.7
Puma8nh 8192 8 70.2 70.2 69.5
Visualizing_soil 8641 4 84.2 84.2 84.2
Delta_elevators 9517 6 75.2 75.2 74.8
Bank-marketing 10578 7 68.2 68.2 67.4
Mozilla4 15545 5 85.8 85.8 85.7
Pulsar-dataset-htru2 17898 8 95.7 95.7 95.7
California 20634 8 70.5 70.5 68.9
California-housing-classi 20640 8 69.8 69.9 68.7
House_8l 22784 8 80.0 80.0 79.3
Bitcoinheist_ransomware 24780 7 66.7 66.7 66.3
Airlines 26969 7 60.6 60.6 60.3
Electricity 38474 8 68.3 68.3 67.4
Click_prediction_small 39926 8 83.2 83.2 83.2
Diabetes130us 71090 7 57.2 57.2 57.0
Run_or_walk_information 88588 6 79.8 79.8 79.4
Wins ($n\leq 250$) 41 28 24
Wins ($2500250<n<$ 2500) 43 34 8
Wins ($n\geq 2500$) 23 20 3
Average rank 1.67 1.79 2.54
Table 3: Size-weighted average up to 16 nodes (SWA16) for optimal and CART. $n$ is the number of samples and $p$ the number of features (before binarization). The results on the data sets with $n\geq 2500$ are shown in detail. The wins and average rank are for all 165 OpenML data sets in our benchmark.
Claim 2.
Optimal methods have a better accuracy-interpretability trade-off than greedy methods (Lin et al., 2020).
To test Claim 2, we compute the size-weighted accuracy (SWA) introduced in Section 5.3 to measure the surface under the accuracy-interpretability Pareto front. Table 3 shows SWA up to 16 nodes for ODTs and CART for all real data sets. ODTs are trained with a maximum depth of four and CART is trained without a depth limit. Since Section 3.4 showed that the pessimistic binomial objective (Binom.) found smaller trees with almost the same accuracy as the accuracy objective, we include it as well in the table. For brevity, we show only the results for the largest data sets (with 2500 or more instances) and summarize the results of the smaller data sets by reporting the number of wins.
These results show that on average the optimal approach achieves a higher SWA than CART. The results also repeat the conclusion of Section 3.4, that training optimal methods with objectives other than accuracy can be advantageous. For small data sets the variance is higher and the wins are more evenly distributed, but for larger data sets, the optimal methods significantly outperform CART, thus verifying Claim 2. For large data sets, the training accuracy is close to the test accuracy for highly regularized models which means that optimal decision trees reliably improve over greedy trees.
5.4.4 Data Efficiency
Claim 3.
The difference between optimal and greedy approaches diminishes with more data (Murthy and Salzberg, 1995; Costa and Pedreira, 2023).
[…]
Figure 16: Results on the synthetic data sets for increasing (a) number of training samples, and (b) number of numeric features.
To test Claim 3, we evaluate ODT and greedy performance on large real data sets and synthetic data with an increasing number of training samples and features.
Fig. 16 shows how ODTs compare with CART for an increasing number of training instances on synthetic data generated from true trees of depth three. For less than 1000 instances, the ODTs are more accurate than both CART with and without a maximum depth limit. For more than 1000 instances, both ODT and CART obtain 100% test accuracy. CART, however, uses 11 leaf nodes to achieve this result, whereas the optimal approach only requires eight (equal to the true tree’s complexity). Both approaches have approximately the same true discovery rate. However, ODT’s false discovery rate is lower. More instances help the ODT method to reduce its false discovery rate.
The depth-constrained CART’s test accuracy plateaus around 1000 training instances at 98%. This shows that CART requires a higher depth limit to obtain the same accuracy as ODTs, regardless of how much data it receives. Even though the true tree depth is three, a maximum depth of four for CART is not enough to recover the tree.
Fig. 16 shows how CART’s accuracy drops when the number of features in the synthetic data increases whereas the optimal approach retains the same perfect accuracy. This difference can be explained by observing the rise in the FDR of CART when the number of features increases together with an increase in the number of leaf nodes: it finds more unnecessary splits. This shows that CART performs worse for an increasing number of features, whereas the optimal approach remains unaffected.
Figure 17: Test accuracies for increasing training samples with and without tuning.
Figure 18: Actual number of leaf nodes when training CART with depth limit 20 for increasing training samples with and without tuning. These numbers far exceed the eight leaf nodes for Optimal in Fig. 17.
Fig. 17 shows the out-of-sample accuracies for increasing training sample sizes on three large real data sets. We train ODTs with a depth limit of 3, CART with a depth limit of 3, and CART without a depth limit. When the methods are not tuned, the optimal approach overfits on small data sets, obtaining a lower test accuracy than depth-limited CART. However, with tuning, this effect disappears and the ODTs’ accuracy is consistently higher than CART’s (depth limited). The difference in performance between tuning and not tuning diminishes for larger training sets. Without a depth limit, CART continues to increase its accuracy for more data. More data does not help depth-limited CART since, at some point, the greedy decisions on what feature and threshold to split on do not change anymore. In those cases, the added data does not lead to different greedy decisions but only makes them more certain.
However, Fig. 18 also shows that CART continues to grow larger trees with up to tens of thousands of nodes. For the Electricity data set, for example, CART and Optimal have roughly similar accuracy for ten thousand training samples. However, the ODT has eight leaf nodes, whereas CART has over a hundred. These results contradict the observation by Oates and Jensen (1997) that greedy tree methods do not perform much better for more data but do yield larger trees with more data. For these data sets, we observe CART performing much better for more data while also resulting in larger trees.
In conclusion, these results refute Claim 3 that the difference between optimal methods and greedy diminished for more data. CART (without a depth limit) can improve performance over ODT with sufficient data. However, unconstrained CART uses unnecessary splits and can result in trees that are orders of magnitude larger than ODTs. Depth-constrained CART may fail to recover an accurate tree even with large training sets and the difference with ODTs does not diminish. Therefore for both depth-constrained and unconstrained CART, we find that they remain different from ODTs with more data.
5.4.5 Model Complexity and Training Data
Claim 4.
The accuracy of greedy trees remains stable when the data size increases linearly with concept complexity (Murthy and Salzberg, 1995).
Figure 19: Results on the synthetic data sets for ground truth trees of increasing complexity and training samples $n=50\cdot 2^{d}$, with $d$ the depth of the ground truth tree.
To test Claim 4, we repeat the experiment by Murthy and Salzberg (1995) by using synthetic data with the number of training samples linear in terms of the number of leaf nodes of the ground truth tree. We set this number to 50 times the number of leaf nodes. Unlike Murthy and Salzberg (1995), we prune the greedy tree and compare it with the optimal tree result instead of comparing it with the ground truth tree.
Fig. 19 shows the effect of linearly increasing the sample size with the true tree complexity. It shows that Greedy’s True Discovery Rate (and False Discovery Rate) decrease (increase) faster than the ODT’s. Interestingly, both methods find trees that have roughly the same number of leaf nodes as the true tree. Regarding Claim 4, both CART with and without a depth limit have a strong decrease in accuracy as the ground truth complexity increases, whereas the ODT’s performance remains close to 100%. Therefore, our results falsify Claim 4. The performance of greedy methods reduces when the true depth increases since increasing true depth requires an increasing number of correct greedy decisions. Since greedy decisions cannot be undone, the probability of making wrong decisions increases.
5.4.6 Overfitting
Claim 5.
Optimal trees are more likely to overfit than greedy trees (Dietterich, 1995).
[…]
Figure 20: Results on the synthetic data for increasing (a) feature noise, and (b) class noise.
We addressed overfitting before when observing the results in Fig. 17. These results showed that without hyperparameter tuning, ODTs are more prone to overfitting than greedy approaches when data is sparse. However, with tuning and with the same size constraint, we observe that ODTs perform better than greedy trees on average.
We further analyze the risk of overfitting by comparing ODTs with greedy on the synthetic data with noise. Fig. 20 shows how both approaches respond to increasing feature and class noise. CART has a similar TDR for increasing feature noise as ODT but a much higher FDR throughout all levels of feature noise. CART with a depth constraint has a lower TDR and accuracy. When feature noise is added, it has a sudden jump in tree size, which corresponds with an increase in the FDR. Each method is only slightly affected by the feature noise.
The impact of class noise is stronger, and we see a steep decrease in accuracy. Depth-limited CART again has a sudden jump in the tree size with a small amount of class noise. As more class noise is added, all methods yield smaller trees and observe a reduced TDR. Only the ODTs observe a growth in the FDR.
These results show that ODTs are not more sensitive to noise than greedy trees. Combined with the previous result on learning trees with a small training sample, this proves that Claim 5 is false when ODTs are properly tuned. At a fixed size, ODTs can overfit more than greedy trees but achieve better performance than greedy at a smaller size. Therefore, the regularization effect of a fixed size on optimal and greedy trees is different and cannot be directly compared.
5.4.7 Expected Tree Depth
Claim 6.
The expected depth of greedy trees remains (in practice) close to the optimal tree depth, even for complex (large) true tree size (Goodman and Smyth, 1988; Murthy and Salzberg, 1995).
Figure 21: The expected depth on the synthetic data is almost always the same.
The expected depth (or question length) is the expected number of branching nodes visited by a test instance. We evaluate Claim 6 about the expected depth of optimal and greedy trees on the synthetic data. Because of the large runtime, we evaluated the synthetic data for true trees of depth 5 on 250 runs rather than the usual 500.
Fig. 21 shows that CART’s expected depth in all cases, for varying training data size, number of features, feature noise, class noise, and ground truth complexity, remains the same as the optimal approach. When CART is depth-constrained, it remains similar to the optimal approach, except when feature noise or some class noise (but not too much) is added. Then, depth-constrained CART yields larger expected depths. Therefore, our results support Claim 6. Although the maximum depth of greedy methods can be significantly larger than the true depth, most deep paths will only be reached by a few samples.
5.5 Scalability of Optimal Decision Trees
A final major difference between optimal and greedy approaches is their scalability. The worst-case runtime of dynamic programming ODT methods grows exponentially with the size of the tree, linearly with the number of samples, and exponentially with the number of binary features. Contrasting this with greedy methods whose runtime only grows linearly with the size of the tree, linearly with the number of features, and log-linearly with the number of samples, it is clear that greedy methods scale better in runtime.
$\operatorname{depth}=2$ $\operatorname{depth}=3$ $\operatorname{depth}=4$
Samples $10^{2}$ $10^{3}$ $10^{4}$ $10^{5}$ $10^{6}$ $10^{2}$ $10^{3}$ $10^{4}$ $10^{5}$ $10^{6}$ $10^{2}$ $10^{3}$ $10^{4}$ $10^{5}$ $10^{6}$
50 features s s s s s s s s s s s s s s m
100 s s s s s s s s s s s s s s m
150 s s s s s s s s s m s s s m m
200 s s s s s s s s s m s s s m -
250 s s s s m s s s s m s m m m -
300 s s s s m s s s m m s m m - -
Table 4: Approximate magnitudes of runtimes for training ODTs using STreeD on synthetic data. s for (sub)seconds, m for minutes, and dashes for runtimes over 2 hours.
Since scalability is one of the limitations of ODTs, we test for what problem sizes the use of ODTs is practically feasible. Table 4 provides an overview of runtimes for the optimal method STreeD in terms of seconds, minutes, and hours. We find that training ODTs up to depth four remains practically feasible for data sets up to approximately 250 binary features for 100,000 instances and 150 binary features for one million instances.
Given such runtime results, and the observation that optimal decision trees are 1) specifically good for training small trees up to depth four (Claim 1), 2) for obtaining trees with a good accuracy-interpretability trade-off (Claim 2), and 3) increasingly so for data sets with many instances, we see the best application of ODTs in finding small interpretable trees for medium to large data sets with relatively few but meaningful and well-prepared features.
5.6 Discussion
We have verified and refuted several claims from the literature about optimal and greedy decision tree methods. Optimal methods outperform greedy approaches when constrained with the same depth limit (up to depth four), verifying Claim 1. They obtain solutions with a better accuracy-interpretability trade-off (verifying Claim 2). Though the greedy trees are larger, the average question length for both methods is roughly the same (verifying Claim 6).
However, size-constrained greedy trees remain worse than ODTs with more data, and unconstrained greedy trees need larger trees for similar ODT performance (refuting Claim 3). For increasing ground truth complexity, the greedy approach requires more data than the optimal approach to retain the same accuracy (refuting Claim 4). Though the variance of the accuracy for small data sets is high, this is true for both methods, and we did not observe more overfitting for the optimal approach (refuting Claim 5). These results solidify the value of optimal trees when size-limited trees are desired.
6 Conclusion
We addressed how to train optimal decision trees (ODTs) and how they compare to greedy approaches and provide recommendations for both (Recommendations 1 and 2).
Since the design of greedy top-down induction (TDI) methods prevents direct accuracy optimization, the literature shows a large variety of learning objectives. We identified and analyzed nine such existing decision tree learning objectives and observed that the concavity of these objectives leads to counter-intuitive results when trained to optimality. Based on this analysis, we propose four novel non-concave objective functions. In contrast to greedy trees, we found that ODTs can optimize out-of-sample accuracy by directly maximizing the in-sample accuracy (provided the data set is large enough), leading to accuracy improvements of more than 1% over objectives such as entropy and Gini impurity. For small data sets, we show that the novel non-concave objectives outperform directly optimizing accuracy because they penalize leaf nodes with close to uniform class distributions more strongly.
We further evaluated four existing and three novel methods for tuning a tree’s complexity across 164 datasets with trees up to depth four, and 131 datasets with trees up to depth five. The experiments show that tuning has a statistically significant impact on the accuracy of at least 1%. The differences between the seven methods are small, specifically for small and large data sets. For medium-sized data sets, tuning the number of nodes, the complexity cost, or the smoothing parameter is best.
Another important contribution of this work is establishing a set of best practices and the supporting analysis framework for comparing ODT and greedy methods. For this, we analyzed previous comparisons and then used the newly established procedure to test six claims from the literature. Our results confirm that (1) ODTs, on average, outperform greedy trees by 1-2% in out-of-sample accuracy for trees trained up to depth four. (2) ODTs, on average, have a better accuracy-interpretability trade-off, which we evaluate with a new metric: the size-weighted accuracy. We refute the claim that (3) differences between greedy and optimal diminish for more data. Depth-constrained greedy methods may fail to recover the true tree, even for large data sets, where ODTs succeed. Unlike previously claimed, (4) greedy methods do not maintain a similar accuracy when the data set size increases linearly with the true tree complexity. ODTs, on the other hand, do. When the complexity of a tree is properly tuned, we find no support for the claim that (5) ODTs are more prone to overfitting than greedy trees. Finally, we find supporting evidence for the claim that (6) the expected depth of greedy trees remains close to the optimal tree.
These results increase our understanding of how to train ODTs and solidify their importance, specifically to obtain small interpretable trees with high accuracy performance.
Future research could investigate the comparison of ODTs with greedy methods without explicit binarization as a preprocessing step. Additionally, the insights from the non-concave objectives could be further exploited, both for ODTs and greedy heuristic design.
Appendix A Data sets
Table LABEL:tab:data_sets lists the OpenML 165 data sets used in the experiments in this paper (Vanschoren et al., 2013; Feurer et al., 2021).
Table 5: List of OpenML data sets used in this paper.
ID Data set Samples Features Binarized features
919 rabe_166 40 1 10
857 bolts 40 7 31
791 diabetes_numeric 43 2 20
729 visualizing_slope 44 3 30
928 rabe_97 46 3 13
777 sleuth_ex1714 47 7 70
817 diggle_table_a1 48 4 40
835 analcatdata_vehicle 48 4 9
942 chscase_health 50 3 27
787 witmer_census_1980 50 4 40
346 aids 50 4 30
476 analcatdata_bankruptcy 50 5 50
874 rabe_131 50 5 50
892 sleuth_case1201 50 6 60
780 rabe_265 51 6 60
713 vineyard 52 2 18
467 analcatdata_japansolvent 52 8 80
790 elusage 55 2 20
864 sleuth_ex2015 60 7 61
887 mbagrade 61 2 11
755 sleuth_ex1605 62 5 50
894 rabe_148 66 5 50
929 rabe_176 70 3 30
804 hutsof99_logis 70 7 37
1015 confidence 72 3 30
893 visualizing_hamster 73 5 50
859 analcatdata_gviolence 74 8 51
945 kidney 76 6 36
459 analcatdata_asbestos 83 3 14
472 lupus 87 3 26
946 visualizing_ethanol 88 2 14
40683 postoperative-patient-dat 88 8 22
1055 cm1_req 89 8 17
891 sleuth_case1202 93 6 44
731 baskball 96 4 40
865 analcatdata_neavote 100 2 10
875 analcatdata_chlamydia 100 3 16
726 fri_c2_100_5 100 5 50
1463 blogger 100 5 13
829 fri_c1_100_5 100 5 50
754 fri_c0_100_5 100 5 50
916 fri_c3_100_5 100 5 50
461 analcatdata_creditscore 100 6 37
45615 appendicitis_test_edsa 106 7 70
771 analcatdata_michiganacc 108 3 24
890 cloud 108 7 61
736 visualizing_environmental 111 3 30
782 rabe_266 120 2 19
448 analcatdata_boxing1 120 3 21
1556 acute-inflammations 120 6 15
714 fruitfly 125 4 24
40681 mux6 128 6 6
924 humandevel 130 1 10
867 visualizing_livestock 130 2 15
1075 datatrieve 130 8 71
885 transplant 131 3 30
921 analcatdata_seropositive 132 3 23
444 analcatdata_boxing2 132 3 21
974 hayes-roth 132 4 11
719 veteran 137 7 36
1013 analcatdata_challenger 138 2 12
784 newton_hema 140 3 30
902 sleuth_case2002 147 6 24
969 iris 150 4 39
955 tae 151 5 30
1026 grub-damage 155 8 48
40669 corral 160 6 6
748 analcatdata_wildcat 163 5 29
747 servo 167 4 19
801 chscase_funds 185 2 16
446 prnn_crabs 200 7 61
733 machine_cpu 209 6 49
796 cpu 209 7 58
895 chscase_geyser1 222 2 20
41538 conference_attendance 246 6 24
464 prnn_synth 250 2 20
730 fri_c1_250_5 250 5 50
744 fri_c3_250_5 250 5 50
911 fri_c2_250_5 250 5 50
776 fri_c0_250_5 250 5 50
1495 qualitative-bankruptcy 250 6 18
811 rmftsa_ctoarrivals 264 2 20
450 analcatdata_lawsuit 264 4 30
1073 jEdit_4.0_4.2 274 8 66
43 haberman 306 3 26
1524 vertebra-column 310 6 60
818 diggle_table_a2 310 8 79
1167 pc1_req 320 8 31
925 visualizing_galaxy 323 4 36
1011 ecoli 336 7 52
1048 jEdit_4.2_4.3 369 8 68
860 vinnie 380 2 13
1025 analcatdata_germangss 400 5 16
900 chscase_census6 400 6 60
906 chscase_census5 400 7 70
907 chscase_census4 400 7 70
908 chscase_census3 400 7 70
909 chscase_census2 400 7 70
1511 wholesale-customers 440 8 65
764 analcatdata_apnea3 450 3 20
814 chscase_vine2 468 2 18
724 analcatdata_vineyard 468 3 29
765 analcatdata_apnea2 475 3 20
767 analcatdata_apnea1 475 3 20
884 fri_c0_500_5 500 5 50
749 fri_c3_500_5 500 5 50
870 fri_c1_500_5 500 5 50
792 fri_c2_500_5 500 5 50
886 no2 500 7 70
750 pm10 500 7 70
335 monks-problems-3 554 6 15
333 monks-problems-1 556 6 15
949 arsenic-female-bladder 559 4 40
334 monks-problems-2 601 6 15
997 balance-scale 625 4 16
770 strikes 625 6 60
795 disclosure_x_tampered 662 3 30
931 disclosure_z 662 3 30
827 disclosure_x_noise 662 3 30
774 disclosure_x_bias 662 3 30
44596 blood-transfusion-service 748 4 34
44593 blood-transfusion-service 748 4 34
44594 blood-transfusion-service 748 4 34
44595 blood-transfusion-service 748 4 34
44597 blood-transfusion-service 748 4 34
1464 blood-transfusion-service 748 4 34
1014 analcatdata_dmft 797 4 20
912 fri_c2_1000_5 1000 5 50
813 fri_c3_1000_5 1000 5 50
743 fri_c1_1000_5 1000 5 50
799 fri_c0_1000_5 1000 5 50
45604 dummy 1000 6 60
43255 1StudentPerfromance 1000 7 43
741 rmftsa_sleepdata 1024 2 14
934 socmob 1156 5 31
1462 banknote-authentication 1372 4 40
991 car 1728 6 21
962 mfeat-morphological 2000 6 43
914 balloon 2001 1 10
772 quake 2178 3 30
40704 Titanic 2201 3 5
737 space_ga 3107 6 60
44127 phoneme 3172 5 49
871 pollen 3848 5 50
728 analcatdata_supreme 4052 7 24
720 abalone 4177 8 73
40983 wilt 4839 5 50
1460 banana 5300 2 20
803 delta_ailerons 7129 5 50
725 bank8FM 8192 8 73
816 puma8NH 8192 8 80
923 visualizing_soil 8641 4 31
819 delta_elevators 9517 6 47
44126 bank-marketing 10578 7 51
1046 mozilla4 15545 5 40
45558 Pulsar-Dataset-HTRU2 17898 8 80
45028 california 20634 8 80
45578 California-Housing-Classification 20640 8 80
843 house_8L 22784 8 75
45037 BitcoinHeist_Ransomware 24780 7 48
42493 airlines 26969 7 67
44156 electricity 38474 8 68
43901 click_prediction_small 39926 8 56
45022 Diabetes130US 71090 7 41
40922 Run_or_walk_information 88588 6 60
Additionally, we perform some experiments on datasets with more than 100,000 samples:
  • covertype (ID 44121) with 566,602 samples, 10 features, and 100 binarized features.
  • Higgs (ID 44129) with 940,160 samples, 24 features, and 240 binarized features.