### v2

# Optimal or Greedy Decision Trees? Revisiting their Objectives, Tuning, and Performance

Abstract
Recently there has been a surge of interest in optimal decision tree (ODT) methods that globally optimize accuracy directly, in contrast to traditional approaches that locally optimize an impurity or information metric. However, the value of optimal methods is not well understood yet, as the literature provides conflicting results, with some demonstrating superior out-of-sample performance of ODTs over greedy approaches, while others show the opposite. Through a novel extensive experimental study, we provide new insights into the design and behavior of learning decision trees. In particular, we identify and analyze two relatively unexplored aspects of ODTs: the objective function used in training trees, and tuning techniques. Thus, we address these three questions: what objective to optimize in ODTs; how to tune ODTs; and how do optimal and greedy methods compare? Our experimental evaluation examines 11 objective functions, six tuning methods, and six claims from the literature on optimal and greedy methods on 180 real and synthetic data sets. Through our analysis, both conceptually and experimentally, we show the effect of (non-)concave objectives in greedy and optimal approaches; we highlight the importance of proper tuning of ODTs; support and refute several claims from the literature; provide clear recommendations for researchers and practitioners on the usage of greedy and optimal methods; and code for future comparisons.
Keywords: optimal decision trees, CART, objectives, complexity tuning, classification
1 Introduction
Decision trees (DTs) are among the most-used (interpretable) machine learning (ML) models. Despite their simplicity, they can learn complex non-linear relationships in data and their human comprehensibility answers the need for interpretable models in high-stake domains (Rudin, 2019; Arrieta et al., 2020), provided the trees are small. Optimal decision trees (ODTs) specifically, which provably optimize an objective for a given size limit, provide small but accurate models on many tabular data sets and thus combine high performance with interpretability (Piltaver et al., 2016; Loh, 2014; Carrizosa et al., 2021).
Because training optimal decision trees with respect to a size limit is NP-hard (Hyafil and Rivest, 1976), most early decision tree learning methods were greedy top-down induction heuristics. Such methods, like CART (Breiman et al., 1984) and C4.5 (Quinlan, 1993), locally optimize some impurity or information gain metric for each branching node.
Consequently, greedy decision tree learning has been extensively studied given its long history; important examples being splitting criteria (Mingers, 1989b; Buntine and Niblett, 1992; Shih, 1999; Raileanu and Stoffel, 2004; Wang and Xia, 2017) and pruning techniques to avoid overfitting (Mingers, 1989a; Esposito et al., 1997; Patil et al., 2010).
In contrast, optimal decision tree research is a much younger field, with the last decade seeing major advancements. The main topic of research has been improving scalability by reducing runtimes and supporting larger data sets. Researchers have employed a variety of techniques such as (mixed) integer programming (MIP) (Bertsimas and Dunn, 2017; Verwer and Zhang, 2017), constraint programming (Verhaeghe et al., 2020), Boolean satisfiability (Hu et al., 2020; Narodytska et al., 2018) and dynamic programming with bounds (Aglin et al., 2020a; Lin et al., 2020; Demirović et al., 2022). Whereas ten years ago optimal methods were limited to small data sets, due to algorithmic advancements and an increase in computation power, we can now use a recent dynamic programming approach (van der Linden et al., 2023) to analyze data sets with up to hundreds of thousands of samples.
Given that the main technical innovations for optimal methods are relatively recent, unlike in the field of greedy heuristics, principled ways of using optimal decision trees for out-of-sample performance have been comparatively under-explored. Whereas the field of greedy decision trees shows a large variety of splitting criteria, pruning methods, and tuning approaches, optimal decisions are almost exclusively trained by maximizing accuracy, possibly additionally penalizing the number of nodes (the sparse objective). Tuning, if done at all, is performed in different ad hoc manners, e.g., tuning the number of nodes or depth of the tree. Practices differ from paper to paper, which hinders direct comparisons.
Moreover, early comparisons between optimal and greedy approaches were limited in scope and contained claims and hypotheses that we can now refute (Section 5). Murthy and Salzberg (1995) lacked a scalable ODT method and therefore confined their analysis on synthetic data. Bertsimas and Dunn (2017) trained ODTs using MIP, but lack of scalability constrained most of their analysis to data sets of only 100 instances or trees with a maximum depth of two. For larger problems, their approach did not converge to optimality; therefore, the support for several of their claims remained uncertain.
Though these and other studies (Lin et al., 2020; Demirović et al., 2022) report an average improvement of the out-of-sample performance versus greedy heuristics, others have criticized ODTs for overfitting (Dietterich, 1995), observed worse results for ODTs compared to greedy heuristics (Zharmagambetov et al., 2021; Marton et al., 2024), and questioned the adjective ‘optimal’ (Sullivan et al., 2024). These contradictory findings illustrate the need for a more thorough understanding of the concept of optimal decision trees.
This motivated us to conduct a thorough experimental evaluation of existing decision tree methods, both greedy and optimal, focusing on objective functions used during training and different tuning approaches. Based on our findings, we motivate new objective functions and tuning approaches that are specific to optimal methods that globally optimize the objective. To support our study, we conducted the largest evaluation to this date concerning optimal and greedy decision tree methods, taking into account 11 different objective functions, six tuning approaches, 180 real-world and synthetic data sets (small and large), and trees that go beyond small tree-depth limits. This provides us with a wealth of data to analyze and improves our understanding of how to apply greedy and optimal approaches for training decision trees.
From our new insights obtained on training ODTs, we also discuss the implications for decision tree learning in general. To keep the scope of this study manageable, we chose to limit this study to axis-aligned binary classification trees with hard splits, which are arguably the most common type of decision trees. In more detail, we contribute the following:
  • In Section 3, we analyze and experimentally compare nine existing greedy decision tree accuracy objectives. Since we observe that the strict concavity of these objectives, as required by greedy top-down inducting approaches (Kearns and Mansour, 1996), is counterproductive when trained to optimality, we also introduce and experiment with two new non-concave objectives. Our experiments show that greedy and optimal approaches respond oppositely to the (non-)concavity of these objectives. Additionally, we show the benefit of objectives that include a regularizing component for noisy data (in addition to the regular tree size tuning).
  • In Section 4, we compare six complexity tuning approaches for ODTs, four of which were proposed before, and two new tuning approaches that we introduce here. Our experiments highlight the importance of tuning optimal decision tree methods, and that (surprisingly) the accuracy differences between the commonly used tuning methods are small, although there are differences in resulting tree size and runtimes.
  • In Section 5, we analyze previous comparisons between greedy and optimal approaches, formulate best practices for future comparisons, and provide data and code to support proper benchmarking. [^1] The code will be made public on paper publication. We apply these practices in evaluating six claims from the literature on the performance of greedy and optimal trees:
    1. Claim 1:
      Optimal methods under the same depth constraint (up to depth four) find trees with 1-2% higher out-of-sample accuracy than greedy methods (Bertsimas and Dunn, 2017; Verwer and Zhang, 2017; Demirović et al., 2022).
    2. Claim 2:
      Optimal methods obtain a better accuracy-interpretability trade-off than greedy methods (Lin et al., 2020).
    3. Claim 3:
      The difference between optimal and greedy approaches diminishes with more data (Murthy and Salzberg, 1995; Costa and Pedreira, 2023).
    4. Claim 4:
      The accuracy of greedy trees remains stable when the data size increases linearly with concept complexity (Murthy and Salzberg, 1995).
    5. Claim 5:
      Optimal trees are more likely to overfit than greedy trees (Dietterich, 1995).
    6. Claim 6:
      The question length of greedy trees remains (in practice) close to that of optimal trees (Goodman and Smyth, 1988; Murthy and Salzberg, 1995).
    Our results support Claims 12, and 6, and refute Claims 34, and 5.
The remainder of the paper is organized as follows. The next section provides a general overview of the field of optimal decision trees. Sections 3, 4, and 5 are mostly self-contained sections, each dedicated to a single major research question as outlined above, each with its corresponding related work, technical details, experiments, and conclusions. Section 6 draws an overarching conclusion.
2 Related Work
This section introduces previous literature on decision tree learning with a focus on optimal methods. For longer reviews, we refer to surveys by Safavian and Landgrebe (1991); Kotsiantis (2013); Costa and Pedreira (2023) and Blockeel et al. (2023).
Decision tree learning started several decades ago with AID (Morgan and Sonquist, 1963), a recursive approach to regression analysis, later adapted for classification in CHAID (Kass, 1980). Since then, two of the most popular decision tree learning algorithms have been CART (Breiman et al., 1984) and ID3 (Quinlan, 1986) with its successor C4.5 (Quinlan, 1993). Each of these uses top-down induction (TDI) to greedily partition the data by finding a split that is locally optimal according to an information or impurity criterion. Overfitting is prevented either by early stopping rules such as a minimum information gain or by post-pruning the tree. We discuss splitting criteria in Section 3.1.
Besides TDI heuristics, other heuristics include stochastic gradient descent (Norouzi et al., 2015), coordinate descent (Carreira-Perpinán and Tavallali, 2018; Dunn, 2018; Bertsimas and Dunn, 2019), evolutionary algorithms (Barros et al., 2011; Guidotti et al., 2024), swarm optimization (Panhalkar and Doye, 2022), and look-ahead (Kiossou et al., 2024). These metaheuristics typically obtain better trees by considering a larger search space than TDI heuristics but do not guarantee to find the globally optimal tree.
Because computing optimal trees is NP-hard (Hyafil and Rivest, 1976; Cox et al., 1989; Murphy and McCraw, 1991), historically most approaches have been heuristics. Although a few early optimal dynamic programming (Schumacher and Sevcik, 1976; Payne and Meisel, 1977; Miyakawa, 1985; Cox et al., 1989; Nijssen and Fromont, 2007, 2010) and an extreme-point tabu search (Bennett and Blue, 1996) approaches were proposed, only recently, with increased compute and algorithmic advancements, interest in optimal trees resurged.
This resurge started with mixed-integer programming (MIP) formulations for ODTs (Bertsimas and Dunn, 2017; Verwer and Zhang, 2017) with several consecutive improvements (Verwer and Zhang, 2019; Zhu et al., 2020; Günlük et al., 2021; Hua et al., 2022; Alès et al., 2024; Aghaei et al., 2024; Liu et al., 2024). The advantages of these MIP methods are that they can find splits with arbitrary thresholds on the continuous features and can easily be adapted by adding linear constraints or changing the objective, including objectives that operate on the whole tree instead of summing the objectives of independent leaf nodes. The disadvantage is poor scalability because of a weak linear relaxation and the inability of the MIP solver to recognize the independence between subtrees.
Around the same time (maximum) satisfiability (SAT) formulations were proposed (Narodytska et al., 2018; Hu et al., 2020; Janota and Morgado, 2020; Avellaneda, 2020; Shati et al., 2023; Alòs et al., 2023). These SAT models focus on finding perfect trees of minimum size and MaxSAT is used to maximize the training accuracy for a fixed size limit. Similarly, Verhaeghe et al. (2020) find perfect trees of minimum size using constraint programming. They improve performance by exploiting the subtree independence, by caching and reusing solutions to subproblems, and by pruning the search through bounds.
Similar techniques are exploited in dynamic-programming (DP) based approaches, which a recent survey (Costa and Pedreira, 2023) indicates as the most promising approach in terms of scalability. DP exploits the independent subtree structure and reuses partial solutions to repeated subproblems. The addition of bound-based pruning (Aglin et al., 2020a, b), improved lower bounds (Hu et al., 2019; Lin et al., 2020; Demirović et al., 2022), and a faster subprocedure for trees of depth two (Demirović et al., 2022) have greatly improved the scalability of the basic DP approach. The advantages of DP are the good scalability for realistic use cases, specifically with respect to the number of instances. The disadvantages are the need for binarization and an exponential runtime with respect to the number of features and maximum tree size.
Recent developments for ODTs are incorporating continuous features (Mazumder et al., 2022; Shati et al., 2023; Brita et al., 2025); exploring the Rashomon set of all close to optimal models (Xin et al., 2022; Semenova et al., 2023); quantifying and reducing explanation redundancy (Izza et al., 2022; Audemard et al., 2022); improving anytime performance (Kiossou et al., 2022; Demirović et al., 2023); improving memory usage (Aglin et al., 2022); and applying ODTs to other objectives such as regression (Zhang et al., 2023; van den Bos et al., 2024), quantile regression (Lemaire et al., 2024), fairness constraints (Aghaei et al., 2019; van der Linden et al., 2022; Jo et al., 2023), robustness (Vos and Verwer, 2022; Justin et al., 2022), survival analysis (Zhang et al., 2024; Huisman et al., 2024), prescriptive policy generation (Bertsimas et al., 2019; Jo et al., 2021; van der Linden et al., 2023), and learning MDP policies (Vos and Verwer, 2023).
Alternatively, others have studied decision trees with soft (probabilistic) decision splits, also known as randomized trees (Blanquero et al., 2021, 2022), or trees with oblique (multivariate) splits (Bertsimas and Dunn, 2017; Zhu et al., 2020; Blanquero et al., 2021; Boutilier et al., 2023; Engür and Soylu, 2024). However, such models are less human-comprehensible and out-of-scope for this paper.
In summary, the recent literature shows a surge in methods for and applications of ODTs. Advances in scalability make it now possible to do a more in-depth analysis of how ODTs should be trained and how they compare to the traditional greedy approaches.
3 The Optimization Objective for Optimal Decision Trees
Decision tree learning objectives typically optimize two parts: some accuracy objective (i.e., accuracy or one of its proxies, such as information gain) and a tree-complexity objective (e.g., number of nodes). In this section, we focus on the first: the accuracy objective. In Section 4, we discuss the tree-complexity objective.
Section 3.1 explains why existing greedy splitting criteria do not optimize accuracy directly. We transform these greedy criteria to ODT objectives and observe in our analysis of these objectives in Section 3.2, that the strict concavity traditionally required by greedy heuristics (Kearns and Mansour, 1996) is not helpful when training ODTs. Therefore, we introduce two novel non-concave objectives in Section 3.3. We then empirically compare all these objectives in Section 3.4 and discuss our findings in Section 3.5. Our main finding is that greedy learners perform best with strictly concave splitting criteria, whereas optimal learners achieve the best performance with non-strictly-concave objectives.
3.1 Greedy Proxies for Accuracy
Although the goal of decision tree learning is to maximize accuracy, TDI methods rarely optimize accuracy directly but instead optimize a proxy, such as the reduction in the $\chi^{2}$-statistic in CHAID (Kass, 1980), Gini impurity by CART (Breiman et al., 1984), and information gain (entropy) by ID3 (Quinlan, 1986) and C4.5 (Quinlan, 1993).
Figure 1: (Left) Three splitting heuristics compared. The horizontal axis shows the binary class distribution expressed as the probability of the first class, and the vertical axis shows the corresponding splitting criterion value (lower is better). (Right) Geometric interpretation of the weighted mean error of two children when $p$, $p_{1}$, and $p_{2}$ represent the class distributions of the parent and the two children respectively. The length of the arrow indicates the improvement in the splitting criterion value. Adapted from Flach (2012).
The reason why TDI methods do not optimize accuracy directly is that an accuracy splitting criterion is often unable to find an improving split in unbalanced data. When splitting a node, TDI methods evaluate all possible splits and choose the split that minimizes the splitting criterion value for the resulting class distributions among the new nodes of each possible split. Fig. 1 visually explains why using accuracy as a splitting criterion is worse at distinguishing improving splits than Gini impurity or entropy. The left side shows the function values for accuracy, Gini impurity, and entropy for binary classification. The right side shows how a locally optimal split can be found geometrically. When splitting a node with a probability of the first class of $p$ into two new nodes with probabilities $p_{1}$ and $p_{2}$, the new weighted splitting criterion value can be found by drawing a straight line from the criterion value at point $p_{1}$ to $p_{2}$. The intersection of the straight line at $p$ is the sum of the weighted criterion value of the two nodes. For Gini impurity and entropy, this value is always lower than the criterion value of the parent node, because both functions are strictly concave. Accuracy, however, is not strictly concave, and when $p\leq 0.5,p_{1}\leq 0.5$, and $p_{2}\leq 0.5$ (or equivalently, all are greater than or equal to $0.5$), the weighted sum of the criterion value of the child nodes is the same as that of the parent node. Moreover, for any values $p_{1}\leq 0.5$ and $p_{2}\leq 0.5$ the weighted sum of the criterion values is the same, and therefore no distinction can be made between these splits. Thus TDI heuristics require strictly concave splitting criteria (Kearns and Mansour, 1996) and therefore do not optimize accuracy directly.
Figure 2: Objective values for different objective functions for a single leaf node. (Left) The leaf node size is fixed at $n=100$. (Right) The misclassifications are fixed at $e=20$. Surprisingly, the value of the strictly concave objectives increases for a fixed error and increasing leaf node size.
3.2 Analysis of ODT Objectives
To increase our understanding of greedy and optimal decision tree learning approaches, we analyze accuracy and eight other existing greedy splitting and pruning criteria and rewrite them as ODT objectives: Gini impurity, square root Gini, entropy, minimum error, binomial pessimistic error (Binom.), minimum description length (MDL, two encodings: Quinlan and Mehta), and Bayesian. In Appendix A, we rewrite each of these as a function $f$ with as input the number of instances $n$ that reach this leaf and the number of misclassifications $e$ in this leaf. Let $L$ be a set of leaf nodes of size $n$ and with $e$ misclassifications, then the minimization objective of the whole tree becomes $\sum_{(n,e)\in L}f(n,e)$.
The left of Fig. 2 shows the values of these objectives when the leaf node size is fixed but the number of misclassifications increases. The accuracy is a straight line since every misclassification is counted equally. Both the pessimistic binomial score and the minimum error follow the accuracy tightly, with only a small additional cost for higher misclassifications. All other objectives follow roughly the same pattern: the first misclassifications in a node are penalized most and the additional penalty for extra misclassifications decreases.
Similarly, on the right of Fig. 2, when the number of misclassifications in a node is fixed but the leaf node size is changed, the accuracy is a straight line. The pessimistic binomial score and the minimum error again follow accuracy closely. Interestingly, for all other objectives, the objective value increases when the node size increases. Since lower values are preferred, this means these objectives penalize larger leaf nodes more than smaller leaf nodes with the same misclassifications.
It is counter-intuitive that the objective increases for larger nodes with a fixed error. Table 1 shows some examples of relative objective values that are unexpected. For example, according to the entropy criterion, it is better to have two nodes of size 4, with two misclassifications in the first node and zero in the second, than one node of size 8 with one misclassification. Entropy strongly values a pure node, even if this means a higher misclassification rate. Other objectives, such as MDL, value two nodes of size four and two with two misclassifications in the first and none in the second, more than one node of size six with also two misclassifications. Again, a small pure node is valued, even though the node of size four with two misclassifications has a high probability of being misclassified, for example, in the presence of class noise.
Table 1: Pure nodes are overvalued, resulting in splits with pure nodes (e.g., $(n,e)=(4,0)$) and nodes that are labeled randomly $(4,2)$, rather than keeping one node with the same misclassifications $(8,2)$ or even less in case of entropy $(8,1)$.
Objective Expected (lower is better) Observed
Gini impurity $f(8,2)\leq f(4,2)+f(4,0)$ $f(8,2)=3.000$, $f(4,2)+f(4,0)=2.000$
Entropy $f(8,1)\leq f(4,2)+f(4,0)$ $f(8,1)=2.174$, $f(4,2)+f(4,0)=2.000$
MDL (Quinlan) $f(6,2)\leq f(4,2)+f(2,0)$ $f(6,2)=4.708$, $f(4,2)+f(2,0)=4.377$
MDL (Mehta) $f(6,2)\leq f(4,2)+f(2,0)$ $f(6,2)=5.513$, $f(4,2)+f(2,0)=5.409$
Bayes $f(6,2)\leq f(4,2)+f(2,0)$ $f(6,2)=4.379$, $f(4,2)+f(2,0)=4.321$
3.3 Novel Non-Concave Objectives
The odd behavior of the greedy criteria in Fig. 2 and Table 1 is a result of their strict concavity. (Strict) concavity is not a requirement for ODTs because ODTs do not consider splitting criteria and can search beyond a non-improving split. Therefore, we here introduce two non-concave objectives by Noel et al. (2023) that have not previously been used in decision tree learning.
M-loss:
The first is the M-loss, here rewritten in terms of $n$ and $e$:
$f(n,e)=n\left(\frac{1}{1-\frac{e}{n}}-1\right)\,.$
L-loss:
The second objective function that they propose is called the L-loss. Rewritten in terms of $n$ and $e$ this becomes
$f(n,e)=n\Bigg{(}\frac{1}{\sqrt{1-\left(\frac{e}{n}\right)^{2}}}-1\Bigg{)}\,.$
Fig. 3 shows the values these new functions take. For easy comparison, accuracy and Gini impurity are also included in the plot. In contrast to the strictly concave functions, the left side of Fig. 3 shows how the first misclassifications in a leaf node are penalized less, whereas nodes with a (close to) balanced class distribution are heavily penalized. The right side shows that increasing the leaf node size while keeping the number of misclassifications constant, decreases the penalization. Therefore, we hypothesize that these objectives obtain the desired property to penalize nodes with a close-to-equal class distribution more strongly.
3.4 Experiments
In our experiments we aim to answer the following questions:
  1. 1.
    What is the difference between the objectives on out-of-sample accuracy when trained to optimality on the training data?
  2. 2.
    What difference can be observed between the objectives when trained to optimality on the training data or when greedily optimized using TDI heuristics?
  3. 3.
    How do different objectives respond to noise and data set size?
Figure 3: The new objectives show opposite behavior to the strictly concave objectives. Left, the leaf node size is fixed at $n=100$. Right, the misclassifications are fixed at $e=20$.
3.4.1 Experiment Setup
We empirically compare the objectives on a large benchmark set from OpenML (Vanschoren et al., 2013; Feurer et al., 2021). [^2] The code will be made public on paper publication. For the sake of scalability, we selected all binary classification data sets with 50 or fewer features, of which eight or fewer numeric features and no large text features, with no missing values, with at most 100,000 instances, and at least 40 instances. We take the most recent version of the data set and omit duplicates or data sets that only differ in the random seed, resulting in 180 data sets. We split each data set into five folds, creating five train and test pairs each consisting of four and one fold respectively. We list all data sets used in Appendix E and also include a histogram of the data set sizes.
We implemented all the ODT objectives in the ODT method STreeD (van der Linden et al., 2023) because of its scalability and flexibility in supporting new objectives. [^3] https://github.com/algtudelft/pystreed STreeD is a DP approach that requires binary features. Therefore, we binarize the numeric training data with thresholds on the ten quantiles, and the categorical data with one-hot encoding (with at most ten categories). The test data is binarized in the same way. We experimented with other binarization approaches, but noted no significant impact with regard to the analysis presented here. See Section 5.2 for the impact of binarization on CART. See Appendix D.2 for a brief evaluation of the impact of binarization on ODTs.
Additionally, we evaluate on synthetic data sets where we can control the amount of noise. We follow the synthetic data setup from Murthy and Salzberg (1995); Bertsimas and Dunn (2017) and Dunn (2018). We generate $n$ random training instances with $p$ numeric features, uniformly distributed over $[0,1]$. For a given noise strength $f$, we add feature noise by adding noise uniformly drawn from $[-f,f]^{p}$. Again, we binarize the numeric features by threshold predicates on 10 quantiles per numeric feature. We generate a random binary tree on this binarized data of a maximum depth $d$ with at most $2^{d}$ leaf nodes. We choose random splits on the data such that each leaf node contains at least 5 instances. The binary labels of each leaf node are assigned alternately, such that no split leads to two leaf nodes with the same label. After this, we add class noise to a given percentage $c$ of the data by flipping its label. For each training set, we create a corresponding test set without noise of $1000$ instances per leaf node in the generated tree.
While keeping the other values constant, we test with changing the amount of feature noise ($f=0$, $0.2$, $0.4$, $0.6$, $0.8$, $1.0$) and the amount of class noise ($c=0\%$, $10\%$, $20\%$, $30\%$, $40\%$, $50\%$), while also changing the number of instances ($n=50$, $100$, $1000$). We repeat each configuration 1000 times and report averages over these 1000 runs.
We train ODTs up to depth four while tuning the number of branching nodes using cross-validation. [^4] A depth-three tree has at most eight leaf nodes and seven branching nodes and a depth-four tree has at most 16 leaf nodes and 15 branching nodes. In cross-validation, for data sets with up to 100 instances, we use 20 folds; for up to 250 instances, we use ten folds; and otherwise, we use five.
Figure 4: Comparing ODT objectives for $\operatorname{max-depth}=4$. (a) Orange (blue) indicates (non-)concave. The average accuracy and number of leaf nodes over all data sets and folds are shown, sorted by the average rank. (b) Nemenyi critical distance rank test. The average rank per objective is plotted and objectives with a rank difference smaller than the critical distance (CD) at p-value $0.05$ are grouped by a black bar.
We test the objectives on TDI heuristics in our own implementation of CART. We use cost-complexity tuning with accuracy as the pruning objective. For comparison with ODT, we train CART using a depth limit of four on the same binarized data sets. In Section 5.2, we evaluate the impact of these choices.
The average rank is our main performance metric: for each data set split, we round the test accuracy to one decimal and then rank all methods. If multiple methods have the same accuracy, they are all assigned the average rank. E.g., if two methods have the same best score, they both get rank $1.5$. We then report the mean rank over all data sets.
3.4.2 Optimal Decision Tree Results
Fig. 4 shows the average rank, test accuracy, and the number of leaf nodes per objective for trees of depth four on the OpenML benchmark. On average, the best ranking objectives with the best test accuracy are accuracy and its slight variations minimum error and the binomial pessimistic error, and the novel non-concave functions M-loss and L-loss. Maximizing training accuracy is ranked second, although the difference from the top objective is not significant. The pessimistic binomial objective achieves the best test accuracy with a lower average number of leaf nodes than the accuracy objective.
However, the results are close. Fig. 4b shows the results of a Nemenyi critical distance rank test to test the significance (Demšar, 2006). This test computes the critical distance (CD) between the average ranks of two methods to be statistically significant. Fig. 4b shows that all the non-concave objectives are not significantly different for depth four. All the non-concave objectives are significantly better than all the concave objectives.
The runtimes of optimizing the objectives are close: the objective with the lowest runtime (L-loss) is on average 1.9 times faster (geometric mean) than the slowest objective (Bayes).
In Appendix D.1, we show training accuracy results for a selection of data sets for three objectives (accuracy, Gini impurity, and M-loss) for an increasing number of nodes.
3.4.3 Greedy Heuristics Results
Figure 5: Comparing greedy objectives for $\operatorname{max-depth}=4$. The strictly concave objectives (orange) significantly outperform the non-concave objectives (blue).
For comparison, we also train greedy trees with the objectives listed above. Fig. 5 shows the resulting out-of-sample performance. In comparison with the ODTs, the ranking of the objectives is almost reversed, which is in line with our analysis in Section 3.2, that the TDI approach requires strictly concave objectives. The Nemenyi critical distance rank test in Fig. 5b shows that all strictly concave objectives (except square-root Gini) are significantly better than all non-concave objectives for TDI heuristics. The smaller tree sizes of the non-concave objectives show that the greedy heuristic gets stuck early with these objectives and finds no improving splits. Optimizing Gini impurity yields 1.2% higher out-of-sample accuracy than directly optimizing the in-sample accuracy. These results confirm that the traditional Gini and entropy are among the top choices.
3.4.4 Noisy Synthetic Data
To further investigate the performance of the ODT objectives, we compare their relative performance on synthetic data for varying data set sizes and levels of feature and class noise. We limit the comparison to the top four objectives and include Gini impurity as the top representative of the concave objectives.
Fig. 6 shows that for the smallest data sets ($n=50$) with an increasing level of feature noise, the binomial pessimistic and the minimum error obtain (statistically) significantly better test accuracy ($+1\%$) than the other objectives. Gini impurity performs slightly worse. With more data ($n=100,1000$), the advantage of the binomial pessimistic and minimum error disappears, whereas Gini impurity performs relatively even worse, up to $2\%$.
Fig. 6 shows that the differences for increasing amounts of class noise are smaller. However, Gini impurity still performs worse than the other objectives, especially for larger data sets ($n=1000$) and more class noise. When the class noise reaches $50\%$, half of the labels are flipped and the training labels essentially are completely random. That is why the relative test accuracy for $50\%$ class noise goes back to zero.
[…]
Figure 6: Relative test accuracy of ODT objectives compared to optimizing training accuracy directly. The concave Gini impurity performs significantly worse. For small datasets with feature noise, the binomial pessimistic and minimum error perform significantly better (the error bar represents the 95% confidence interval).
3.5 Discussion
Previous work has extensively compared greedy splitting criteria, which we extend to optimal decision trees. Our results show that optimizing accuracy directly is a good choice for ODTs, specifically when the number of training samples increases because, with sufficient training data, the training accuracy closely approximates the test accuracy. This shows that the strict concavity of objectives such as Gini impurity and entropy, is not an inherently necessary or desirable property, but a limitation imposed by the greedy TDI approach. Optimal and greedy training procedures respond differently to the objectives and therefore best practices of one approach do not necessarily translate to the other. This opens the question of how the performance of non-strictly-concave objectives for small data sets could be exploited in greedy heuristics.
We hypothesized that non-concave objectives may perform better than accuracy for noisy data. Although we did not measure a significant difference for the new objectives M-loss and L-loss, we found two objectives that can outperform optimizing accuracy directly for small noisy data sets: the pessimistic binomial and minimum error objective. Both of these objectives include a regularizing component, which –in addition to the regularization effect of tuning the tree size– helps to prevent overfitting. This also results in slightly smaller models, while retaining the same average accuracy.
4 Tuning the Complexity of Optimal Decision Trees
Most decision tree learning approaches make a trade-off between training accuracy and model complexity to prevent overfitting. For ODTs, several complexity tuning methods have been used but without an in-depth empirical comparison. Therefore, this section analyses the effect of complexity tuning methods for optimal decision trees, starting with an overview of complexity tuning methods. We find that optimal decision trees perform significantly better with tuning than without, but that existing tuning techniques perform similarly.
4.1 Tuning Approaches
Currently, ODT approaches mostly tune the following hyperparameters: tree depth (Aglin et al., 2020a), tree size (Demirović et al., 2022), complexity cost (Lin et al., 2020), and the minimum support (Nijssen and Fromont, 2007). In Appendix B, we provide additional information on these approaches. On top of the existing ODT tuning approaches, we evaluate two new approaches:
Question length:
[…]
Smoothing:
Because of the good performance of the minimum error objective in Section 3.1, we generalize this approach by tuning the Laplace smoothing parameter. The Laplace smoothing approach (Flach, 2012) assumes in a leaf node that for each class, $x$ extra instances exist. With $|\mathcal{K}|$ the number of classes, the accuracy objective becomes
$f(n,e)=\frac{n(e+x)}{n+|\mathcal{K}|x}\,.$
4.2 Experiments
The experiments aim to answer the following questions: how do the ODT tuning methods compare in out-of-sample accuracy and how do they respond to increasing noise?
[…]
Figure 7: Complexity tuning results for ODTs of (a) $\operatorname{max-depth}=4$ and (b) $\operatorname{max-depth}=5$ for five runs on (a) 180 data sets and (b) 157 data sets.
4.2.1 Experiment setup
We evaluate each tuning method on the same 180 OpenML and synthetic data sets described in Section 3.4.1. Regardless of the tuning method, we impose a maximum depth constraint of four or five. For each tuning method, we select at most $k$ different parameter settings. For depth, these settings are selected using equal linear spacing between the $k$ options. For the others, the $k$ settings are selected using equal spacing in the log scale. In the results presented here, we set $k=16$ (because a depth-four tree can have zero to fifteen branching nodes, i.e., 16 options). We provide more details on how the values are chosen and experiments for other values of $k$ in Appendix B. All tuning methods are implemented in STreeD (van der Linden et al., 2023).
For most data sets, we obtain trees well within the maximum depth limit of five, and therefore we do not extend this analysis to larger depth limits.
4.2.2 Results on Real Data Sets
Figs. 7 and 7 show the performance of the complexity tuning method on the OpenML data sets. We exclude data sets from the analysis if any method exceeded the two-hour time-out: for depth four, no data sets are removed, and for depth five, 23 data sets. Surprisingly, the results show that all tuning methods obtain similar accuracies: there is no statistically significant difference between any of the approaches. In terms of optimizing accuracy, the only conclusion is that using any of the tuning approaches is better than no tuning.
However, other differences can be observed between the methods. For example, tuning the question length, minimum support, or depth yields slightly larger trees. Furthermore, in Appendix B, we measure the runtime and test the performance of the methods for different numbers of parameter settings $k$. This shows that tuning only the depth of the tree is by far the fastest approach.
[…]
Figure 8: Relative test accuracy of ODT tuning methods compared to tuning the size (number of nodes) of the tree. No tuning, tuning the minimum support, and tuning the question length are significantly worse than the other tuning approaches when noise is present.
Inspecting the results shows that the differences among tuning methods are largest for medium-sized data sets (several hundred to ten thousand samples). For example, on the 82 data sets with 250 or more and 10,000 or fewer samples, a Wilcoxon signed rank test shows tuning the size is statistically significantly better than tuning the depth (p-value $\approx 0.01$).
4.2.3 Results on Synthetic Data
To further test the tuning methods’ performance, we evaluate them on the synthetic data sets. Fig. 8 shows that as the amount of feature noise increases, tuning becomes increasingly more important. It also shows that tuning the question length is slightly worse than the other tuning methods. No clear differences can be observed between the other tuning methods.
Fig. 8 shows larger differences when increasing the amount of class noise. Again, not tuning is significantly worse when noise is present. Tuning the minimum support and the question length is significantly worse than the other tuning methods when class noise is present. At $50\%$ class noise, this difference obviously disappears because then all training labels are basically decided randomly.
Interestingly, Fig. 8 shows no significant difference in tuning only the depth versus the more expensive procedures of tuning the size, the cost-complexity, or the smoothing level.
4.3 Conclusion
In conclusion, complexity tuning of ODTs is necessary. On our real data sets, all previously used tuning approaches, obtain similar accuracy results. On the noisy synthetic data, on the other hand, tuning the question length or minimum support is significantly worse than the alternatives. Tuning the depth is more time efficient than other approaches, obtains the same accuracy, but does so with slightly more nodes. One new tuning approach is promising: tuning the amount of smoothing. Based on our experiments in Sections 3 and 4, we recommend training ODTs as follows:
{mdframed}
Recommendations 1 (Training Optimal Decision Trees).
  1. 1.
    Optimize the same loss at train and test time for optimal decision trees.
    Optimizing the accuracy on average yields the best out-of-sample accuracy. Do not optimize concave proxies such as Gini impurity or entropy.
  2. 2.
    In noisy data sets, consider an objective with an additional regularizing effect.
    In noisy data sets, optimizing objectives such as the pessimistic binomial and minimum error may perform better than maximizing accuracy as they have an additional regularization effect by encouraging model sparsity.
  3. 3.
    Tune the complexity of optimal decision trees.
    Training ODTs with hyperparameter tuning is significantly better than training without hyperparameter tuning.
  4. 4.
    Tune the size, complexity cost, or smoothing parameter; or, in the case of large data sets, the depth.
    Tuning the size, complexity cost, smoothing, or depth, on average, yields similar out-of-sample results. However, tuning the depth yields larger trees. For large data sets, tuning is less important, and tuning the depth is more runtime efficient.
5 Comparing Optimal and Greedy Decision Trees
To understand the differences between greedy and optimal approaches for learning decision trees, we collected claims from the literature and evaluated these claims with extensive experiments. Below, we list the claims discussed in this section and summarize our results:
Claim 1:
Optimal methods under the same depth constraint (up to depth four) find trees with 1-2% higher out-of-sample accuracy than greedy methods (Bertsimas and Dunn, 2017; Verwer and Zhang, 2017; Demirović et al., 2022).
Supported:
We evaluate the accuracy of depth three and four trees on 180 data sets and find an average improvement of 1.3% and 1.0% of optimal over greedy approaches.
Claim 2:
Optimal methods obtain a better accuracy-interpretability trade-off than greedy methods (Lin et al., 2020).
Supported:
We evaluate the accuracy of trees with 1 to 16 nodes on 180 data sets and find that the size-weighted accuracy of optimal methods is, on average, significantly higher than that of greedy methods. The size-weighted accuracy is a new metric we propose in Section 5.3 to measure the accuracy-interpretability trade-off.
Claim 3:
The difference between optimal and greedy approaches diminishes with more data (Murthy and Salzberg, 1995; Costa and Pedreira, 2023).
Refuted:
Experiments on synthetic and real data show that size-constrained greedy trees do not improve after some point and stay worse than optimal. In contrast, the performance of size-unconstrained greedy trees keeps improving with more samples, while also growing much larger trees than ODTs.
Claim 4:
The accuracy of greedy trees remains stable when the data size increases linearly with concept complexity (Murthy and Salzberg, 1995).
Refuted:
On synthetic data generated from a random decision tree, the performance of optimal decision trees remains stable when the random tree’s size is increased, and the performance of greedy trees diminishes.
Claim 5:
Optimal trees are more likely to overfit than greedy trees (Dietterich, 1995).
Refuted:
With hyperparameter tuning, we do not find significant performance differences between optimal and greedy methods with small numbers of samples (up to 250) nor more sensitivity to noise.
Claim 6:
The question length of greedy trees remains (in practice) close to the optimal tree depth (Goodman and Smyth, 1988; Murthy and Salzberg, 1995).
Supported:
Our experiments on synthetic data show that the question length of optimal and greedy methods remains similar in (almost) all (practical) scenarios. Only for very noisy data, does CART yield much longer question lengths.
In this section, we first review previous greedy-optimal comparisons, which we use to design a set of best practices for future comparisons. The rest of the section details the experiments we performed to evaluate each of the claims from existing literature. In Section 5.5 we discuss the scalability of ODTs.
5.1 Previous Comparisons
Year Author(s) Method     Greedy with same size limit     Greedy without size limit     Small and large data sets     Beyond small trees     Optimal tuned (correctly)     Greedy tuned (correctly)
Papers that propose ODT methods
2007 Nijssen and Fromont DL8
2017 Bertsimas and Dunn OCT
2019 Verwer and Zhang BinOCT
2020 Lin et al. GOSDT
Hu et al. MaxSAT
2021 Günlük et al. ODT
2022 Demirović et al. MurTree
Hua et al. RS-OCT
Mazumder et al. Quant-BnB
2024 Liu et al. BNP-OCT
Alès et al. CTT
2025 Brita et al. ConTree
Other papers that compare ODTs with greedy DTs
1995 Murthy and Salzberg -
2021 Zharmagambetov et al. -
2024 Sullivan et al. MAPTree
Marton et al. GradTree
Table 2: Simplified overview of the comparisons between greedy and optimal methods in the literature. Ideally, a comparison checks all columns. 1) Compare methods under the same size constraint; 2) compare (greedy) methods without a size constraint; 3) compare on small and large data sets (> 10.000 instances); 4) compare optimal methods beyond depth three; 5) tune optimal methods (correctly); and 6) tune greedy methods (correctly).
In Appendix C, we list previous comparisons between greedy and optimal approaches. These comparisons can be grouped into ODT papers that propose new ODT methods, and other papers. From these comparisons, the following general observations can be made about how greedy and optimal methods are compared:
  • Both greedy and optimal methods are often not correctly tuned, or not tuned at all. When comparing with CART, many papers show several modifications to how CART is trained and tuned. We assess the impact of each of these modifications on the accuracy in Section 5.2 below. Additionally, several papers evaluated ODTs without tuning the complexity. As shown in Section 4.2, this significantly worsens the ODT performance. Others evaluate ODTs with a substantial restriction on the tree size, resulting in shallow underfitting trees.
  • Almost all ODT papers compare with CART under the same tree-size constraint and draw a positive conclusion on the ODTs’ performance. The other papers typically compare ODTs under a size constraint with unconstrained non-optimal approaches and sometimes draw negative conclusions about the ODTs’ performance. Both comparisons are useful for different purposes. One simply evaluates out-of-sample performance, while the other evaluates the claim that ODTs have a better accuracy-interpretability trade-off. We further discuss measuring this trade-off in Section 5.3 below.
  • Several comparisons are limited to small data sets and small trees (e.g., less than 10000 instances or trees of maximum depth three). This is typically because of scalability limitations. The improved scalability of optimal methods allows us to analyze larger data sets and trees.
Table 2 summarizes our observations about previous comparisons. We recommend that future comparisons between ODTs and greedy approaches should check all the columns in this table. We summarize our recommendations as follows:
{mdframed}
Recommendations 2 (Greedy-Optimal Comparisons).
  1. 1.
    Compare both with and without constraining the sizes of the decision trees.
    Optimal decision trees optimize performance for a particular size; therefore, greedy methods should be equally constrained in size to compare fairly.
  2. 2.
    Compare performance on both small and large data sets.
    While experiments on small data are more efficient to run, their results often do not carry over to larger data sets.
  3. 3.
    Evaluate both small and large trees.
    Previous comparisons often compare trees of depth two. This optimization problem is too simplistic, and the results do not always carry to a larger depth.
  4. 4.
    Tune both greedy and optimal methods and ensure an equal comparison.
    Comparisons between greedy and optimal trees at a fixed depth are unequal since the methods respond differently to size constraints. If one wants to compare trees up to depth four, for example, both approaches should be tuned up to a maximum depth of four, not trained with a fixed depth of four.
5.2 Training CART
In the comparisons reviewed in Appendix C and summarized above, ODTs are often compared to a modified version of CART, for example, to allow for a direct comparison under similar circumstances. We test the impact of these modifications to assess the validity of these previous comparisons and inform future comparisons. We assess the following typical modifications: 1) tuning the depth instead of the complexity cost; 2) binarizing the feature data; or 3) running CART while imposing an additional depth constraint.
We compare CART’s performance with these modifications against unmodified CART on the 180 OpenML data sets used before. We approximate the unconstrained CART with a maximum depth of 20. We set the constrained depth limit to four, because of its common use in ODT comparisons. As before, the binarized data has up to ten binary features per continuous feature by using thresholds on ten quantiles or one-hot encoding of categorical features with a maximum of ten categories.
In addition to the depth limit, we apply cost-complexity pruning as done in RPart (Therneau et al., 2023): we train a fully expanded tree and obtain the cost-complexity path with all possible cost-complexity values from that tree and use the geometric mean to get the midpoints of those values. We use cross-validation to find the best cost-complexity parameter among the midpoints and retrain a tree on the full training data with this parameter.
Unlike RPart, we take the best performing cost-complexity parameter, and not the largest complexity cost which performs within one standard error of the best performing one. In our preliminary tests, this resulted in larger trees but better out-of-sample accuracy.
Figure 9: Results for CART (Gini) when trained with(out) binarization, with(out) a depth limit, and when using depth or cost-complexity tuning. Using a depth limit (as indicated in blue) significantly impacts performance. Tuning the depth has a more negative impact for large maximum depths. Binarization with 10 quantile thresholds has no significant impact on the accuracy but does impact the tree size.
Fig. 9 shows CART’s performance under these modifications. The largest differences are between the depth-constrained and the unlimited depth variant. Binarization has only a small impact on the performance (this does not necessarily generalize for more coarse binarizations). When a strict depth limit is imposed, tuning the depth instead of the complexity cost has a small impact but for the unlimited depth case, this significantly hurts CART’s performance. Fig. 9 also shows significantly different tree sizes for CART’s modifications. Both binarization and depth tuning result in larger trees. From these results, we can conclude the following best practices:
{mdframed}
Recommendations 3 (Training CART).
  1. 1.
    Training CART with a depth limit should be clearly stated.
    CART trained with a depth limit results in significantly different results than CART without a depth limit.
  2. 2.
    Tuning the depth of CART instead of the complexity cost should be avoided.
    Tuning CART’s depth rather than the cost-complexity yields larger trees.
  3. 3.
    Training CART on binarized data should be clearly stated.
    Depending on the binarization, training on binarized data may or may not significantly harm the performance.
5.3 Accuracy-Interpretability Trade-Off
ODT papers typically compare ODTs with greedy heuristics under a similar size constraint. This is partly motivated by the definition of ODTs because ODTs are defined as trees that maximize training performance under a given size limit. No theoretical claim is made about its out-of-sample performance or the performance without a size constraint. In fact, finding an optimal tree without a size constraint is trivial: it is obtained by splitting in any way until no further split can be made.
Figure 10: Typical accuracy-interpretability trade-off for untuned greedy and optimal decision trees. ODTs have a slight advantage for small size limits but both methods converge for large size limits.
It is also motivated by the claim that ODTs have a better accuracy-interpretability trade-off. Since one of the oft-cited motivations for decision trees is their comprehensibility and large trees with hundreds of nodes can hardly be called human-comprehensible (Piltaver et al., 2016), evaluating an algorithm’s ability to obtain small performant trees is important.
Typically, the tree size and test accuracy are plotted against each other, as done in Fig. 10, to assess this trade-off. Such plots show the relative performance for different size limits and also the saturation point: when adding more nodes gives no improvement in accuracy. Optimal methods typically reach this saturation point earlier and greedy methods eventually catch up by obtaining the same accuracy but with larger trees.
Such comparisons require one figure per data set. To enable easier comparisons across a large number of data sets, we propose a new metric: the size-weighted accuracy (SWA). The purpose of this metric is to express the accuracy-interpretability trade-off with one number that represents the ‘surface’ under the accuracy-interpretability Pareto front seen in Fig. 10. Since we are mostly interested in the performance of small trees, we set the weight of the accuracies obtained for a tree with $i$ leaf nodes to $1/i$. We define the size-weighted accuracy ($\operatorname{SWA}_{n}$) as the weighted average of the obtained trees of maximum size $n$:
$\operatorname{SWA}_{n}=\frac{1}{\sum_{i=1}^{n}\frac{1}{i}}\sum_{i=1}^{n}\frac{\operatorname{acc}_{i}}{i}\,.$ (1)
Since not every algorithm can directly set the number of leaf nodes, we propose to run the algorithm with different complexity parameters (e.g., the complexity-cost parameter $\lambda$) and record the resulting number of leaf nodes and test accuracy for each run. If multiple runs yield the same number of leaf nodes, then average these test accuracies. For missing tree sizes, linearly interpolate test accuracies using results from the nearest smaller and larger tree sizes. If the largest tree size obtained is less than $n$, assign the test accuracy of this largest tree to all larger sizes up to $n$. If larger trees result in smaller test accuracy, replace it with the larger value for smaller trees, since we measure the Pareto front.
For example, for the synthetic data in Fig. 10, we computed trees up to 20 leaf nodes. CART obtains a $\operatorname{SWA}_{20}$ of $84.0\%$, whereas the optimal approach obtains a $\operatorname{SWA}_{20}$ of $84.7\%$.
5.4 Experiments on Literature Claims
To evaluate the claims made in previous papers, we compare optimal and greedy methods on synthetic and real data sets, using the best practices introduced in Recommendations 1-3.
5.4.1 Experiment Setup
We again evaluate on both the OpenML and synthetic data sets introduced in Section 3.4.1. We add two more real data sets to evaluate Claim 3: covertype and Higgs. These two were chosen for their large number of samples, allowing us to investigate performance under various sample sizes (566,602 and 940,160 respectively). For the synthetic data, unless otherwise specified, we set the true tree depth to three, the number of instances $n=1000$, the number of numeric features $p=3$, the feature noise $f=0$, and the class noise $c=0\%$. We test with changing the number of instances ($n=50$, $100$, $250$, $1000$, $10000$), the number of features ($p=2$, $4$, $6$, $8$), and with changing the noise as done in Section 3.4.1. Additionally, we add synthetic data sets with a linear separator instead of a tree as the ground truth. The weights of the linear separator are chosen randomly from a normal distribution. We repeat each configuration 500 times and report averages over these 500 runs.
We evaluate CART and ODT on the binarized data to eliminate this difference between the two methods and focus only on the difference between greedy and optimal search. Comparing on binarized data is commonly done (Lin et al., 2020; Demirović et al., 2022). Section 5.2 shows that the impact of this binarization on CART is small for the data sets in our benchmark. In Appendix D.2, we also evaluate the effect of binarization for ODTs.
In our synthetic tree experiments, apart from the familiar test accuracy and number of leaf nodes, we also measure the following:
True Discovery Rate (TDR):
The TDR is the percentage of splits in the ground truth tree that are recovered in the trained tree (higher is better).
False Discovery Rate (FDR):
The FDR is the percentage of the splits in the trained tree that are not part of the ground truth tree (lower is better).
Question Length:
The question length is the average number of branching nodes an instance visits when evaluated (lower is better).
5.4.2 Out-of-Sample Accuracy
Claim 1.
Optimal methods under the same depth constraint (up to depth four) find trees with 1-2% higher out-of-sample accuracy than greedy methods (Bertsimas and Dunn, 2017; Verwer and Zhang, 2017; Demirović et al., 2022).
Figure 11: Out-of-sample accuracy of CART and ODTs compared on five runs for 180 data sets. (a) Optimal (blue) versus CART (red). CART without a depth limit performs best but yields much larger trees. (b) Nemenyi critical distance rank test for optimal versus CART. The average rank per method is plotted and methods with a rank difference smaller than the critical distance (CD) at p-value $0.05$ are grouped by a black bar. With the same depth limit, optimal performs significantly better than CART, confirming Claim 1.
Figure 12: The difference between the optimal and CART out-of-sample accuracy for both small (left) and larger (right) data sets. The dashed line indicates the average difference.
To evaluate Claim 1, we evaluate both the ODT approach and CART on the OpenML data sets with a depth limit of three and four. We also compare with CART without a depth limit. Fig. 11 shows that CART (without a depth limit) performs better (but not significantly) than the ODT approach with a depth-four depth limit, but CART yields much larger trees.
When compared under the same depth constraint, as stated in Claim 1, optimal significantly outperforms greedy with an average improvement of 1.3% and 1.0% for depths three and four respectively. Since optimal algorithms optimize the complete decision tree, instead of greedily improving the tree, they can achieve better scores.
Fig. 12 shows the distribution of the differences between optimal and greedy when trained with maximum depth four. For small data sets ($n\leq 250$), the average advantage of the optimal approach is $0.2\%\pm 0.4$ (mean $\pm$ standard error). Some large accuracy differences occur because the test sets for these small data sets are so small that a single misclassified instance can increase or decrease accuracy by several percent. For larger data sets ($n>250$), the average improvement of optimal over CART is $1.6\%\pm 0.2$. The difference is larger and the standard error is lower. These results confirm Claim 1.
5.4.3 Accuracy-interpretability trade-off
Claim 2.
Optimal methods obtain a better accuracy-interpretability trade-off than greedy methods (Lin et al., 2020).
To test Claim 2, we compute the size-weighted accuracy (SWA) introduced in Section 5.3 to measure the surface under the accuracy-interpretability Pareto front. We train ODTs with the top four non-concave and top two concave objectives from Section 3 and CART with the traditional Gini impurity objective. ODTs are trained with a maximum depth of four and CART is trained without a depth limit.
Figure 13: Comparing the test size-weighted average (SWA) of several ODT objectives and CART (see Section 5.3). (a) The non-concave objectives are blue, the concave objectives are orange, and CART is green. (b) The ODT minimum error, accuracy, and M-loss are significantly better than the other approaches. All ODT approaches are significantly better than CART, confirming Claim 2.
Fig. 13 shows that on average the optimal approach achieves a higher SWA than CART, thus verifying Claim 2. For large data sets, the training accuracy is close to the test accuracy for highly regularized models which means that optimal decision trees reliably improve over greedy trees. This result also repeats the conclusion of Section 3.4, that non-strictly-concave objectives are to be preferred over concave objectives for optimal methods.
5.4.4 Data Efficiency
Claim 3.
The difference between optimal and greedy approaches diminishes with more data (Murthy and Salzberg, 1995; Costa and Pedreira, 2023).
To test Claim 3, we evaluate ODT and greedy performance on large real data sets and synthetic data with an increasing number of training samples and features.
Fig. 14 shows how ODTs compare with CART for an increasing number of training instances on synthetic data generated from ground-truth trees of depth three. For less than 1000 instances, the ODTs are more accurate than both CART with and without a maximum depth limit. For more than 1000 instances, both ODT and CART obtain 100% test accuracy. CART, however, uses 11 leaf nodes to achieve this result, whereas the optimal approach only requires eight (equal to the true tree’s complexity). Both approaches have approximately the same true discovery rate. However, ODT’s false discovery rate is lower. More instances help the ODT method to reduce its false discovery rate.
[…]
Figure 14: Results on the synthetic tree data sets for increasing (a) number of training samples, and (b) number of numeric features; and (c) on the synthetic linear data sets for increasing number of training samples and numeric features, to evaluate Claim 3.
The depth-constrained CART’s test accuracy plateaus around 1000 training instances at 98%. This shows that CART requires a higher depth limit to obtain the same accuracy as ODTs, regardless of how much data it receives. Even though the true tree depth is three, a maximum depth of four for CART is not enough to recover the tree.
Fig. 14 shows how CART’s accuracy drops when the number of features in the synthetic data increases whereas the optimal approach retains the same perfect accuracy. This difference can be explained by observing the rise in the FDR of CART when the number of features increases together with an increase in the number of leaf nodes: it finds more unnecessary splits. This shows that CART performs worse for an increasing number of features, whereas the optimal approach remains unaffected.
Fig. 14 additionally shows that when learning from synthetic data with a linear separator as the ground truth, CART without a depth constraint achieves higher accuracy than ODT, but with more data availability, CART also generates much larger trees with only a small gain in accuracy. Increasing the number of features in the synthetic data makes the classification function harder to learn. The relative accuracy performance of the methods stays roughly the same, but CART requires many more nodes. In all cases, ODTs perform better than CART with the same depth constraint.
Figure 15: Test accuracies for increasing training samples with and without tuning.
Figure 16: Number of leaf nodes for CART for increasing training samples with and without tuning. These numbers far exceed the eight leaf nodes for ODTs in Fig. 15.
Fig. 15 shows the out-of-sample accuracies for increasing training sample sizes on three large real data sets. We train ODTs with a depth limit of three, and CART with and without a depth limit of three. When the methods are not tuned, the optimal approach overfits on small data sets, obtaining a lower test accuracy than depth-limited CART. However, with tuning, this effect disappears and the ODTs’ accuracy is consistently higher than CART’s (depth limited). The difference in performance between tuning and not tuning diminishes for larger training sets. Without a depth limit, CART continues to increase its accuracy for more data. More data does not help depth-limited CART since, at some point, the greedy decisions on what feature and threshold to split on do not change anymore. In those cases, the added data does not lead to different greedy decisions but only makes them more certain.
However, Fig. 16 also shows that CART continues to grow larger trees with up to tens of thousands of nodes. For the Electricity data set, for example, CART and ODTs have roughly similar accuracy for ten thousand training samples. However, the ODT has eight leaf nodes, whereas CART has over a hundred. These results contradict the observation by Oates and Jensen (1997) that greedy tree methods do not perform much better for more data but do yield larger trees with more data. For these data sets, we observe CART performing much better for more data while also resulting in larger trees.
In conclusion, these results refute Claim 3 that the difference between optimal methods and greedy diminished for more data. CART (without a depth limit) can improve performance over ODT with sufficient data. However, unconstrained CART uses unnecessary splits and can result in trees that are orders of magnitude larger than ODTs. Depth-constrained CART may fail to recover an accurate tree even with large training sets and the difference with ODTs does not diminish. Therefore, for both depth-constrained and unconstrained CART, we find that they remain different from ODTs with more data.
5.4.5 Model Complexity and Training Data
Claim 4.
The accuracy of greedy trees remains stable when the data size increases linearly with concept complexity (Murthy and Salzberg, 1995).
Figure 17: Testing Claim 4 on the synthetic data sets for ground truth trees of increasing complexity and training samples $n=50\cdot 2^{d}$, with $d$ the depth of the ground truth tree.
To test Claim 4, we repeat the experiment by Murthy and Salzberg (1995) by using synthetic data with the number of training samples linear in terms of the number of leaf nodes of the ground truth tree. We set this number to 50 times the number of leaf nodes. Unlike Murthy and Salzberg (1995), we prune the greedy tree and compare it with the optimal tree result instead of comparing it with the ground truth tree.
Fig. 17 shows the effect of linearly increasing the sample size with the true tree complexity. It shows that Greedy’s True Discovery Rate (and False Discovery Rate) decrease (increase) faster than the ODT’s. Interestingly, both methods find trees that have roughly the same number of leaf nodes as the true tree. Regarding Claim 4, both CART with and without a depth limit have a strong decrease in accuracy as the ground truth complexity increases, whereas the ODT’s performance remains close to 100%. Therefore, our results falsify Claim 4. The performance of greedy methods reduces when the true depth increases since increasing true depth requires an increasing number of correct greedy decisions. Since greedy decisions cannot be undone, the probability of making wrong decisions increases.
5.4.6 Overfitting
Claim 5.
Optimal trees are more likely to overfit than greedy trees (Dietterich, 1995).
[…]
Figure 18: Testing Claim 5 on the synthetic tree data for increasing (a) feature noise, and (b) class noise; and (c) on the synthetic linear data for increasing feature and class noise.
We addressed overfitting before when observing the results in Fig. 15. These results showed that without hyperparameter tuning, ODTs are more prone to overfitting than greedy approaches when data is sparse. However, with tuning and with the same size constraint, we observe that ODTs perform better than greedy trees on average. We further analyze the risk of overfitting by comparing ODTs with greedy on the synthetic data with noise.
Fig. 18 shows how both approaches respond to increasing feature noise. For all amounts of feature noise, ODT obtains both a higher test accuracy and smaller trees than both CART approaches. The TDR is mostly similar, except for large amount of feature noise which causes the unconstrained CART to train larger trees which results in a higher TDR, but not in higher accuracy. ODT’s FDR is consistently better than CART’s.
Fig. 18 shows that for increasing amounts of class noise, ODT’s test accuracy is again consistently higher than both CART approaches. For large amounts of class noise, unconstrained CART obtains a lower test accuracy than both other approaches and also yields significantly larger trees.
For the synthetic linear data, Fig. 18 shows similar results. In both cases, with little noise, unconstrained CART achieves a higher accuracy but with a much larger tree. However, when either type of noise increases, ODT’s test accuracy becomes higher.
These results show that ODTs are not more sensitive to noise than greedy trees. Combined with the previous result on learning trees with a small training sample, this proves that Claim 5 is false when ODTs are properly tuned. Without tuning, given a fixed size limit, ODTs can overfit more than greedy trees. However, ODTs achieve the same performance as greedy trees at smaller size limits, which means there is no difference in overfitting after tuning. Therefore, the regularization strength of a fixed size on optimal and greedy trees is different and cannot be directly compared.
5.4.7 Question Length
Claim 6.
The question length of greedy trees remains (in practice) close to the optimal tree depth (Goodman and Smyth, 1988; Murthy and Salzberg, 1995).
Figure 19: The question length on the synthetic data is almost always the same, except for data sets with a strong noise component, mostly supporting Claim 6.
The question length is the expected number of branching nodes visited by a test instance. We evaluate Claim 6 about the question length of greedy trees on the synthetic data.
Fig. 19 shows that the question length of depth-constrained CART remains similar to the optimal approach. Also without a depth limit, CART’s question length in almost all cases –for varying training data size, number of features, and ground truth complexity– remains the same as the optimal approach. However, with much feature or class noise, CART’s question length is (much) larger than ODT’s. This difference is most pronounced for more than 30% class noise. Because Murthy and Salzberg (1995) tested with class noise up to 25% this effect was not previously observed.
Theoretically it can be shown that CART’s question length can be arbitrarily larger than optimal (Garey and Graham, 1974). But Claim 6 concerns practical use cases. Therefore, since the difference between CART and ODTs only arises for large noise, we consider Claim 6 supported by our experiments.
5.5 Scalability of Optimal Decision Trees
A final major difference between optimal and greedy approaches is their scalability. The worst-case runtime of dynamic programming ODT methods grows exponentially with the size of the tree, linearly with the number of samples, and exponentially with the number of binary features. Contrasting this with greedy methods whose runtime only grows linearly with the size of the tree, linearly with the number of features, and log-linearly with the number of samples, it is clear that greedy methods scale better in runtime.
Since scalability is one of the limitations of ODTs, we test for what problem sizes the use of ODTs is practically feasible. Table 3 provides an overview of runtimes for the optimal method STreeD in terms of seconds, minutes, and hours when trained on synthetically generated data from a random decision tree. We find that training ODTs up to depth four remains practically feasible for data sets up to approximately 250 binary features for 100,000 instances and 150 binary features for one million instances.
We recollect the observation made before that ODTs are specifically good
  1. 1.
    for training small trees up to depth four (Claim 1);
  2. 2.
    for obtaining trees with a good accuracy-interpretability trade-off (Claim 2);
  3. 3.
    and increasingly so for data sets with many instances (Claim 3).
Therefore, we conclude that the best application of ODTs is finding small interpretable trees for medium to large data sets with relatively few but meaningful and well-prepared features.
$\operatorname{depth}=2$ $\operatorname{depth}=3$ $\operatorname{depth}=4$
Samples $10^{2}$ $10^{3}$ $10^{4}$ $10^{5}$ $10^{6}$ $10^{2}$ $10^{3}$ $10^{4}$ $10^{5}$ $10^{6}$ $10^{2}$ $10^{3}$ $10^{4}$ $10^{5}$ $10^{6}$
50 features s s s s s s s s s s s s s s m
100 s s s s s s s s s s s s s s m
150 s s s s s s s s s m s s s m m
200 s s s s s s s s s m s s s m -
250 s s s s m s s s s m s m m m -
300 s s s s m s s s m m s m m - -
Table 3: Approximate magnitudes of runtimes for training ODTs using STreeD on synthetic data. s for (sub)seconds, m for minutes, and dashes for runtimes over two hours.
6 Conclusion
We experimentally evaluated many different approaches on how to train optimal decision trees (including several novel ones) and how they compare to greedy approaches. Based on the obtained insights, we provide recommendations for training and evaluating optimal and greedy methods (Recommendations 1-3).
Since the design of greedy top-down induction (TDI) methods prevents direct accuracy optimization, the literature shows a large variety of learning objectives. We identified and analyzed nine such existing decision tree learning objectives and observed that the concavity of these objectives leads to counter-intuitive results. Based on this analysis, we evaluate two new non-concave objective functions. By optimizing ODTs for in-sample accuracy, we found significant improvements in out-of-sample accuracy over objectives such as entropy and Gini impurity. For noisy data sets, we show that objectives that have an additional regularizing effect (such as C4.5’s pessimistic binomial error) may improve out-of-sample performance, while also improving the sparsity of the trees.
We further introduced two novel methods for tuning a tree’s complexity and evaluated these against four existing methods across 180 datasets with trees up to depth four, and 157 datasets with trees up to depth five. The experiments show that tuning has a statistically significant impact on the accuracy. The differences in accuracy between the six methods are small, but they may yield smaller or larger trees. Tuning only the depth is promising if runtime performance is a concern. Tests with synthetic noisy data shows that tuning the question length or the minimum support is significantly worse than the other approaches.
Another important contribution of this work is establishing a set of best practices and the supporting analysis framework for comparing ODT and greedy methods. For this, we analyzed previous comparisons and then used the newly established best experimental practices (Recommendations 2) to test six claims from the literature, leading to the following insights. Our results confirm that (1) ODTs, on average, outperform greedy trees by 1-2% in out-of-sample accuracy for trees trained up to depth four. (2) ODTs, on average, have a better accuracy-interpretability trade-off, which we evaluate with a new metric: the size-weighted accuracy. We refute the claim that (3) differences between greedy and optimal diminish for more data. Depth-constrained greedy methods may fail to recover the true tree, even for large data sets, where ODTs succeed. Unlike previously claimed, (4) greedy methods do not maintain a similar accuracy when the data set size increases linearly with the true tree complexity. ODTs, on the other hand, do. When the complexity of a tree is properly tuned, we find no support for the claim that (5) ODTs are more prone to overfitting than greedy trees. Finally, we find supporting evidence for the claim that (6) the question length of greedy trees remains close to that of the optimal tree.
Although unrestricted greedy trees can outperform depth-limited ODTs in accuracy, they can quickly grow so large that they cannot be directly interpreted anymore. Random forests or neural networks already suffice if accuracy is the only concern. However, ODTs are an ideal candidate if interpretability is required, as they achieve a superior accuracy-interpretability trade-off over (unrestricted) greedy trees.
Future research could investigate the comparison of ODTs with greedy methods without explicit binarization as a preprocessing step.
Appendix A Splitting and Pruning Criteria as ODT Objectives
Greedy top-down induction (TDI) methods typically consider two types of criteria: a local splitting criterion that decides how to split a node and a pruning criterion that decides which nodes to keep during pruning. Some greedy methods employ another criterion for pruning such as the minimum description length (Mehta et al., 1995) or the pessimistic error rate in C4.5 (Quinlan, 1993). In this appendix, we list nine such criteria and rewrite them as ODT objectives.
For brevity, we consider only the binary classification case. For all functions, terms that divide by zero (e.g., $e/n$ when $n=0$) or take the log of zero (e.g., $\log_{2}(e/n)$ when $e=0$) are evaluated as zero. We leave out some splitting criteria, such as the twoing rule in CART, if they are not additive or cannot easily be expressed as an objective function for the leaf.
Accuracy:
The most direct way to optimize out-of-sample accuracy is to optimize the in-sample accuracy by minimizing the misclassification score. The leaf node objective in terms of the leaf node size $n$ and the number of misclassifications $e$ is $f(n,e)=e$. A possible disadvantage of this objective is its ‘flatness’: it cannot distinguish between many different solutions. For example, $f(4,2)+f(6,0)=f(5,1)+f(5,1)$.
Gini impurity:
Commonly used in TDI heuristics is the Gini impurity. Weighted Gini impurity scores are obtained by multiplying the Gini impurity by the number of instances in that leaf node. Let $p_{0}=\frac{e}{n}$ denote the probability of the first class and $p_{1}=\frac{n-e}{n}$ the probability of the second class. Then the objective value is
$f(n,e)=n(1-p_{0}^{2}-p_{1}^{2})\,.$
Square root Gini:
Kearns and Mansour (1996) propose to use the square root of the Gini impurity to improve performance on unbalanced data sets:
$f(n,e)=n\sqrt{1-p_{0}^{2}-p_{1}^{2}}\,.$
Entropy:
The second commonly used TDI criterion is entropy (or information gain). Weighted scores are again obtained by multiplying by the size of the leaf node:
$f(n,e)=-\frac{n}{2}(p_{0}\log_{2}p_{0}+p_{1}\log_{2}p_{1})\,.$
Gini impurity and entropy can be expressed as a parameterization of Tsallis entropy (Tsallis, 1988; Wang and Xia, 2017).
Minimum error:
Niblett (1987) estimates the expected error for nodes by assuming that every class has equal probability. It depends on the number of labels $|\mathcal{K}|$, and the count of the majority label $n_{c}$. In binary classification $|\mathcal{K}|=2$ and $n_{c}=n-e$. Therefore, the expected error is
$n\frac{n-n_{c}+|\mathcal{K}|-1}{n+|\mathcal{K}|}=\frac{n(e+1)}{n+2}\,.$
This is equivalent to Laplace smoothing with a smoothing parameter set to one (add-one smoothing) (Flach, 2012). According to Mingers (1989a), the equal-probability assumption of this approach becomes problematic for a large number of classes.
Pessimistic error:
Quinlan (1987) proposed a pessimistic error rate by computing a bound on the expected error rate, which in effect raises the error rate at every leaf by 0.5:
$f(n,e)=e+\frac{1}{2}\,.$
This method is similar to a complexity cost per node of 0.5. Since we cover complexity costs in the next section, we do not consider the pessimistic error in this section.
Binomial pessimistic error (Binom.):
The commonly used C4.5 method (Quinlan, 1993) uses an advanced form of pessimistic error by considering a leaf with $n$ training instances and $e$ misclassifications as a ‘sample’ from a binomial distribution with an unknown misclassifying probability. Since this probability cannot be computed directly, the upper confidence bound of the posterior distribution of this probability, based on a confidence level $\alpha$, is used as the error rate of the leaf node. The confidence interval width $z_{\alpha}$ is the $z$-value from the normal distribution for confidence level $\alpha$. Let $e^{\prime}=e+\frac{1}{2}$ be the pessimistic error. Then the binomial pessimistic error can be expressed as:
$f(n,e)=\begin{cases}n\cdot\left(1-\operatorname{exp}^{\ln{(\alpha)}/n}\right)&\text{if\leavevmode\nobreak\ }e=0\\ e&\text{if\leavevmode\nobreak\ }e=n\\ \frac{e^{\prime}+\frac{z_{\alpha}^{2}}{2}+\sqrt{z_{\alpha}^{2}\left(e^{\prime}\left(1-\frac{e^{\prime}}{n}\right)+\frac{z_{\alpha}^{2}}{4}\right)}}{n+z_{\alpha}^{2}}\cdot n&\text{otherwise.}\end{cases}$
In our experiments, we use the same default $\alpha=0.25$ as C4.5.
Minimum description length (Quinlan):
The minimum description length (MDL) approach states that the best model can be described with the least amount of bits of information because the description length of a model can directly be linked to the posterior probability of a model (Rissanen, 1978; Li and Vitányi, 2008). In practice, the encoding typically consists of two parts: first the encoding of the model and then the encoding of the data that deviates from the model.
Quinlan and Rivest (1989) observe that the cost of encoding a binary string of length $n$ with $e$ ones and $n-e$ zeros can be computed by first encoding the size $n$ of the string, and then the positions of the $e$ ones, with $b$ representing an upper bound on the number of ones that can occur:
$L(n,e,b)=\ln{(b+1)}+\ln{\binom{n}{e}}\,.$ (2)
Then, for every leaf node with $n$ instances and $e$ misclassifications, encode a bit string that specifies the misclassifications with cost $L(n,e,\left\lfloor{n/2}\right\rfloor)$ for binary classification:
$f(n,e)=L(n,e,\lfloor(n+1)/2\rfloor)\,.$
The encoding of the branching feature and the leaf node label are part of the tree complexity cost which we cover in the next section.
Minimum description length (Mehta):
Mehta et al. (1995) observe a sub-optimal coding length for Eq. (2) when $e$ is close to zero, and therefore propose to use stochastic complexity (Krichevsky and Trofimov, 1981; Rissanen, 1997). For binary classification, their formula can be rewritten to
$f(n,e)=e\ln\frac{n}{e}+(n-e)\ln\frac{n}{n-e}+\frac{1}{2}\ln\frac{n}{2}+\ln\pi\,.$
Bayesian:
Decision trees are also commonly trained using a Bayesian approach (Chipman et al., 1998; Denison et al., 1998). These approaches find the maximum likelihood tree given some priors. We present the objective function used in the recent work by Sullivan et al. (2024). For the binary case, they assume that each leaf node can be represented by a Bernoulli distribution with parameter $\theta\in[0,1]$. They assume $\theta\in\operatorname{Beta}(\rho_{0},\rho_{1})$, the Beta distribution with parameters $\rho_{0},\rho_{1}\in\mathbb{R}^{+}$. The values $\rho_{0}$ and $\rho_{1}$ are hyperparameters, but they fix these values to $\rho_{0}=\rho_{1}=2.5$. The error can then be expressed in terms of the Beta function $B$ as follows:
$f(n,e)=\frac{B(e+\rho_{0},n-e+\rho_{1})}{B(\rho_{0},\rho_{1})}\,.$
Sullivan et al. (2024) also add a cost based on the priors, which can be considered a complexity cost, and therefore we leave it out in the discussion in this section.
Almost all previous ODT methods optimize the accuracy, but some have considered other objectives, such as balanced accuracy (Lin et al., 2020) or F1-score (Demirović and Stuckey, 2021; van der Linden et al., 2023). Nijssen and Fromont (2010) have implemented the pessimistic error objective from C4.5 and a Bayesian objective function, but do not discuss their effect on the out-of-sample performance.
Appendix B ODT Tuning Approaches
This appendix provides further details on the ODT hyperparameter tuning approaches, the choice of hyperparameters when limited to $k$ options, and more experiments.
B.1 Details on Existing Tuning Approaches
We evaluate four existing ODT hyperparameter tuning approaches: depth, size, cost-complexity, and minimum support. Here, we provide some background on each of the tuning methods, and how we select the $k$ options for the hyperparameter. For each approach, we always include a setting that provides no limitation on the tree size. The other settings are derived using linear or log-equal distances between settings. The following provides more details per method:
Depth:
A common metric for tree complexity is its depth. DL8.5 (Aglin et al., 2020a, b), for example, tunes the depth of the tree. A possible disadvantage is that the number of nodes increases exponentially with respect to the maximum depth, thus providing only a coarse control of the tree size.
In our experiments, we tune the depth parameter $d\in\{0,...,\operatorname{max-depth}\}$ with equal linear space between the options and with a $\operatorname{max-depth}$ of either four or five.
Size:
The number of nodes can be tuned directly as a hard constraint. Since binary trees always have one more leaf node than branching nodes, the total number of nodes can be counted by the number of branching or leaf nodes alone. This approach is taken by, e.g., MurTree (Demirović et al., 2022) which directly tunes both the maximum depth and the number of branching nodes.
[…]
Complexity cost:
The most common approach in greedy trees is to penalize the cost of adding a node by a factor $\lambda$. This approach is used in most MIP methods (Bertsimas and Dunn, 2017) and also in the optimal sparse approaches such as GOSDT (Lin et al., 2020). Complexity-cost tuning minimizes $\lambda|\mathcal{D}||L|+\sum_{(n,e)\in L}f(n,e)$, with $L$ the set of leaves, $|\mathcal{D}|$ the size of the data set, and $\lambda$ the complexity-cost parameter.
The minimum change in $\lambda$ that may result in a different tree is $\lambda=\frac{1}{|\mathcal{D}|\operatorname{max-depth}}$ because below this value it is never worth adding nodes to increase accuracy. We set the maximum value to $0.05$ and select $k-1$ options from this range using equal log spacing. The minimum step size between values is set to $\frac{1}{|\mathcal{D}|\operatorname{max-depth}}$. In all cases, we add $\lambda=0$.
Note that Lin et al. (2020) recommend for their method GOSDT to use $\lambda\geq\frac{1}{|\mathcal{D}|}$ for faster training, but this setting can exclude larger trees that are more accurate. Additionally, in their experiments, they aim to acquire trees of at most $n$ leaves by setting $\lambda=\frac{1}{n}$. However, this also filters out many trees that have much less than $n$ leaves, since generally a single leaf node already has an accuracy greater than 0.5 for binary classification, so that even a perfect tree with $n$ leaves given such $\lambda$ would have a lower score than a single leaf node. Therefore, both of these settings are too conservative.
Minimum support:
The minimum support is a hard constraint on the minimum number of instances that in a leaf node. Minimum support is more common in the data mining literature, e.g., Nijssen and Fromont (2007). It is also introduced as a soft constraint (Vilas Boas et al., 2021).
In our experiments, we tune the minimum leaf node size $m$ as a percentage of the data set size $|\mathcal{D}|$. The minimum value for $m$ is $\frac{1}{|\mathcal{D}|}$: precisely one sample should end up in each leaf node. For the maximum value, we compute the frequency of the majority class $|\mathcal{D}_{\operatorname{majority}}|$ and set the maximum value of $m$ to $1-\frac{|\mathcal{D}_{\operatorname{majority}}|}{|\mathcal{D}|}$. Any value higher than this will always result in a single leaf node. We obtain $k$ values from this range using equal log spacing, with a minimum step size of $\frac{1}{|\mathcal{D}|}$.
We leave out selecting a tree structure from a fixed set (Günlük et al., 2021), because this approach was chosen specifically to deal with the scalability limitations of MIP.
B.2 Details on the New Tuning Approaches
We set up the two new tuning approaches introduced in the main text as follows:
Question length:
In our experiments, we select $k-1$ options for $\omega$ using equal log spacing from the interval $[\frac{1}{|\mathcal{D}|\operatorname{max-depth}},0.1]$. Additionally, we always add the option $\omega=0$. We also use $\frac{1}{|\mathcal{D}|\operatorname{max-depth}}$ as the minimum step size.
Smoothing:
In our experiments, we select $k-1$ options for $x$ using equal log spacing from the interval $[\frac{1}{\operatorname{max-depth}},0.05|\mathcal{D}|]$. Additionally, we always add the option $x=0$. We also use $\frac{1}{\operatorname{max-depth}}$ as the minimum step size.
B.3 Extended Tuning Experiments
Here, we report additional experiments on the ODT tuning methods. We test the performance of the methods for other values of $k$, the number of hyperparameter options, and we report the average runtime for each approach. While in the main text, we experimented with $k=16$, here we test for $k=5,10,20$. For runtime comparison, we ran all experiments on an Intel Xeon E5-6248R 3.0GHz with 100GB RAM using one thread.
Figure 20: Performance of all tuning methods for $k=5,10,20$ with $\operatorname{max-depth}=4$ on the 180 OpenML data sets. The color indicates the tuning method. Test accuracy is roughly the same for all methods. Increasing $k$ typically yields smaller trees. As expected, runtime scales roughly linear with $k$.
Fig. 20 shows that the choice of $k$ has no major impact on the out-of-sample performance of the tuning methods. In almost all cases, running with more hyperparameter options results in a lower average number of leaf nodes. As expected, increasing $k$ by a factor two, also increases the runtime for each method by a factor of roughly two.
These results also show that the runtimes of all methods are in the same order of magnitude. The only exception to this is tuning the depth. Therefore, the choice of hyperparameter tuning and the number of parameter options to test mostly depends on how much time one is willing to spend on finding small trees and less important for optimizing accuracy.
Appendix C Previous Comparisons between Optimal and CART
This section reviews previous comparisons between optimal and greedy decision tree learning methods, each different in its experiment design and sometimes in its conclusion. We restrict our review to binary axis-aligned trees. First, we summarize the comparisons from papers that proposed ODT methods. Second, we discuss other papers that compare greedy heuristics and optimal methods. Finally, we discuss the best and poor comparison practices as input for a fair comparison method presented after. We highlight the encountered claims that were mentioned above.
C.1 Comparisons in ODT Papers
Papers that propose new ODT methods typically aim to train a decision tree with a given size constraint that achieves the best out-of-sample performance. Nijssen and Fromont (2007, 2010) compare their optimal DL8 algorithm with J48, an implementation of C4.5. When trained on the same discretized data, without a depth limit, but with the same minimum support constraint, DL8 is significantly better for 9 out of 20 data sets and worse for one, while yielding trees that are 1.5 times larger than J48. However, when J48 is trained without the minimum support constraint and with the non-discretized data, J48 outperforms DL8 on out-of-sample accuracy for most data sets.
Bertsimas and Dunn (2017) develop the optimal MIP method OCT and compare it with CART on real and synthetic data. When CART is constrained to the same depth limit as OCT (up to depth four), they conclude that OCT, on average, has a statistically significant 1-2% better out-of-sample accuracy (Claim 1). The largest difference is observed at depth two. They hypothesize that the smaller difference for depths three and four is the result of OCT not reaching optimality within their time limit. When CART is run with a depth limit of ten, it is negligibly better than OCT at depth four.
The main restriction of their analysis is the scalability of OCT. Because of this, they restrict synthetic data analysis to data sets with only 100 instances and two continuous features. They also experiment with data sets up to 1600 instances, but only on ground truth trees of depth two. When training OCT on the synthetic data, they set the maximum depth to the true depth, which prevents overfitting.
Verwer and Zhang (2019) compare the optimal MIP methods BinOCT, DTIP (Verwer and Zhang, 2017), and OCT with depth-constrained CART on data sets with a few thousand instances. They report results without hyperparameter tuning and observe that the ODTs are significantly better for depths two and three and slightly better for depth four (Claim 1).
Lin et al. (2020) propose GOSDT, an ODT method with a sparsity coefficient. They conclude that GOSDT obtains a better accuracy-interpretability trade-off than other methods, including CART (Claim 2). This is based on an experiment on six small data sets with a coarse binarization applied to both GOSDT and CART. CART is tuned using the maximum depth parameter (from one to six), instead of tuning the complexity-cost as is normally done. They tune GOSDT using complexity-cost tuning without a depth limit.
Demirović et al. (2022) compare their optimal MurTree algorithm and CART on binarized data sets with up to 43500 instances and 1163 binary features. They run MurTree for different depths (from one to four) and number of nodes, and CART for different depths (from one to four), and report the best test accuracy for each method. They too report an average out-of-sample improvement of 1-2% over CART (Claim 1).
The same trend appears in other papers that propose new ODT methods. The aim is to show ODTs’ superior performance under a fixed depth limit. An exception is (Alès et al., 2024), who compare with unconstrained greedy approaches. Results are often shown for fixed hyperparameters (Hu et al., 2020; Mazumder et al., 2022; Liu et al., 2024). Scalability limits the analysis to small data sets (Hu et al., 2020; Günlük et al., 2021; Alès et al., 2024) or larger data sets are run only with a maximum depth of two (Hua et al., 2022).
C.2 Other Comparisons
Papers that do not propose new ODT methods typically have another aim: the best out-of-sample accuracy without imposing depth constraints on the tree.
Murthy and Salzberg (1995) compare the greedy approach to known true (optimal) trees on a synthetic benchmark. They observe that the greedy tree is approximately one standard deviation larger than the true tree size while the question length (what they call expected depth) is very close to optimal, which was also observed by Goodman and Smyth (1988) (Claim 6). However, the maximum depth of greedy trees is on average approximately two times higher than the true depth. When they increase the data set size linearly with the true tree complexity, they observe almost no drop in accuracy for the greedy approach (Claim 4). From this result, Costa and Pedreira (2023) hypothesize that the gap between optimal and greedy approaches diminishes for more data (Claim 3).
Dietterich (1995) concludes from the empirical results by Quinlan and Cameron-Jones (1995) that optimal methods are more prone to overfitting (Claim 5). Exhaustively searching through all possible models may yield smaller models, but is also more prone to finding small patterns that do not represent the ground truth. Therefore Dietterich concludes that it is better to train greedy methods with a model complexity penalty.
Zharmagambetov et al. (2021) compare greedy methods with their local search method TAO (Carreira-Perpinán and Tavallali, 2018) and the optimal methods OCT and GOSDT. They conclude that most methods perform similarly, except TAO, which outperforms the other methods. In many cases, CART outperforms OCT and GOSDT. For CART and TAO, they train greedy trees up to depth 30. For OCT, they use the results reported in (Bertsimas and Dunn, 2017), which go up to depth four. They train GOSDT with a high complexity cost, yielding trees that are on average no larger than 3.4 leaf nodes for any of the data sets.
Sullivan et al. (2024) propose MAPTree, a search algorithm that finds the maximum a posteriori tree. They compare with DL8.5, GOSDT, and CART and conclude that MAPTree outperforms these approaches, which leads them to question the ‘optimality’ of ODTs. They observe that DL8.5 is prone to overfitting, while GOSDT is prone to underfitting, and both are sensitive to hyperparameter tuning, whereas MAPTree is not. However, these results are from averaging the performance per hyperparameter setting over all data sets, rather than tuning the hyperparameters for each individual run. Additionally, they evaluate MAPTree without a depth limit, while other methods (including CART) are run with a depth limit.
Marton et al. (2024) learn axis-aligned trees with gradient descent and compare the results a.o. to CART and DL8.5. GradTree outperforms the other methods for binary classification, while CART performs the best for multi-class classification. The ODT approach DL8.5 performs the second worst in both cases: only the evolutionary approach is worse. They tune each method using random search, except for DL8.5 which they fix to a maximum depth of four. The other methods were typically trained up to depths 7-10.
Appendix D Additional Experiments
Section D.1 provides more details on the train and test accuracy of various objectives. Section D.2 measures the effect of binarization for ODTs.
D.1 Train Accuracy
Figure 21: Train and test accuracy for ODTs with different objectives for increasing size limits for the data sets visualizing_slope ($n$=44), aids ($n$=50), confidence ($n$=72), and lupus ($n$=87). Train accuracy is always close. Test accuracy differences are larger.
Figure 22: Train and test accuracy for ODTs with different objectives for increasing size limits for the data sets servo ($n$=167), haberman ($n$=306), vinnie ($n$=380), and pollen ($n$=3848). For haberman, vinnie, and pollen, the train accuracy of Gini impurity is worse.
Fig. 21 and 22 show the train and test accuracies obtained by training optimal decision trees for three objectives for an increasing node limit for small and medium-sized data sets. All trees are trained with $\operatorname{max-depth}=4$. The training accuracy of the three objectives is typically close. For the larger data sets, Gini impurity obtains a lower training accuracy. Optimizing accuracy or M-loss results in (almost) the same training accuracy.
For several of these data sets, the training accuracy plateaus early. For some (other) data sets, the test accuracy remains stable or decreases when trained with more nodes, indicating that the trees are possibly overfitting. The variance in the test accuracy results is large, making it hard to draw conclusions from these figures.
D.2 Binarization
To test the effect of binarizing features, we experiment with changing the number of binary features per original numeric or categorical feature. For all data sets, we compute optimal trees of maximum depth three or four when creating 2, 5, 10, 15, or 25 binary features per original feature. We then measure the test accuracy for each tree and compute the normalized test accuracy by dividing it by the highest test accuracy obtained for that data set. Fig. 23 shows how on average the test accuracy plateaus when creating more and more binary features per original feature. Our selection of 10 binary features per original feature on average is 0.6% less accurate for $\operatorname{max-depth}=3$ and 0.5% less accurate for $\operatorname{max-depth}=4$ than when using 15 binary features per original feature. Fig. 23 also shows the exponential increase in runtime when the number of binary features is increased.
Figure 23: The normalized accuracy for an increasing number of binary features per original feature. At some point, more binary features add little extra information, while significantly increasing the runtime.
What binarization technique works best is a topic that we leave aside in this paper. Other work studies this, such as McTavish et al. (2022) who use ensembles for binarization, and Piccialli et al. (2024) who use counterfactuals for discretization. Recently, also new ODT methods have been proposed that do not require explicit binarization (Mazumder et al., 2022; Shati et al., 2023; Brita et al., 2025).
In this paper, the precise method of binarization is not a major concern. We assume that the binarization is given, so that all methods work on precisely the same data. In this way, we eliminate this difference between methods and can focus in the comparison on the difference between the greedy and optimal approaches.
Appendix E Data sets
Table LABEL:tab:data_sets lists the 180 OpenML data sets used in the experiments in this paper (Vanschoren et al., 2013; Feurer et al., 2021).
Table 4: List of OpenML data sets used in this paper.
ID Data set Samples Features Binarized features Class imbalance
919 rabe_166 40 1 10 0.53
857 bolts 40 7 31 0.65
938 sleuth_ex1221 42 9 90 0.55
40660 analcatdata_fraud 42 11 19 0.69
791 diabetes_numeric 43 2 20 0.60
729 visualizing_slope 44 3 30 0.61
777 sleuth_ex1714 47 7 70 0.57
835 analcatdata_vehicle 48 4 9 0.56
817 diggle_table_a1 48 4 40 0.52
942 chscase_health 50 3 27 0.52
346 aids 50 4 30 0.50
787 witmer_census_1980 50 4 40 0.52
476 analcatdata_bankruptcy 50 5 50 0.50
892 sleuth_case1201 50 6 60 0.52
713 vineyard 52 2 18 0.54
467 analcatdata_japansolvent 52 8 80 0.52
790 elusage 55 2 20 0.56
864 sleuth_ex2015 60 7 61 0.55
887 mbagrade 61 2 11 0.52
755 sleuth_ex1605 62 5 50 0.50
804 hutsof99_logis 70 7 37 0.51
1015 confidence 72 3 30 0.83
893 visualizing_hamster 73 5 50 0.55
859 analcatdata_gviolence 74 8 51 0.58
945 kidney 76 6 36 0.53
459 analcatdata_asbestos 83 3 14 0.55
472 lupus 87 3 26 0.60
862 sleuth_ex2016 87 10 80 0.52
946 visualizing_ethanol 88 2 14 0.51
40683 postoperative-patient-dat 88 8 22 0.73
1055 cm1_req 89 8 17 0.78
479 analcatdata_cyyoung9302 92 9 67 0.79
891 sleuth_case1202 93 6 44 0.61
731 baskball 96 4 40 0.51
465 analcatdata_cyyoung8092 97 10 78 0.75
865 analcatdata_neavote 100 2 10 0.93
875 analcatdata_chlamydia 100 3 16 0.81
1463 blogger 100 5 13 0.68
754 fri_c0_100_5 100 5 50 0.54
46544 dataset_analcatdata_credi 100 6 36 0.73
965 zoo 101 16 20 0.59
45615 appendicitis_test_edsa 106 7 70 0.80
771 analcatdata_michiganacc 108 3 24 0.56
890 cloud 108 7 61 0.70
736 visualizing_environmental 111 3 30 0.52
448 analcatdata_boxing1 120 3 21 0.65
1556 acute-inflammations 120 6 15 0.51
714 fruitfly 125 4 24 0.61
40681 mux6 128 6 6 0.50
924 humandevel 130 1 10 0.50
867 visualizing_livestock 130 2 15 0.81
1075 datatrieve 130 8 71 0.92
885 transplant 131 3 30 0.63
921 analcatdata_seropositive 132 3 23 0.65
974 hayes-roth 132 4 11 0.61
719 veteran 137 7 36 0.69
1013 analcatdata_challenger 138 2 12 0.93
784 newton_hema 140 3 30 0.50
902 sleuth_case2002 147 6 24 0.53
1006 lymph 148 18 47 0.55
969 iris 150 4 39 0.67
955 tae 151 5 30 0.66
1026 grub-damage 155 8 48 0.68
40669 corral 160 6 6 0.56
748 analcatdata_wildcat 163 5 29 0.71
747 servo 167 4 19 0.77
463 backache 180 31 99 0.86
801 chscase_funds 185 2 16 0.53
941 lowbwt 189 9 37 0.52
1012 flags 194 28 117 0.64
446 prnn_crabs 200 7 61 0.50
733 machine_cpu 209 6 49 0.73
796 cpu 209 7 58 0.75
895 chscase_geyser1 222 2 20 0.60
41538 conference_attendance 246 6 24 0.87
464 prnn_synth 250 2 20 0.50
776 fri_c0_250_5 250 5 50 0.50
1495 qualitative-bankruptcy 250 6 18 0.57
811 rmftsa_ctoarrivals 264 2 20 0.62
450 analcatdata_lawsuit 264 4 30 0.93
336 SPECT 267 22 22 0.79
1073 jEdit_4.0_4.2 274 8 66 0.51
23499 breast-cancer-dropped-mis 277 9 37 0.71
1121 badges2 294 10 43 0.71
40710 cleve 303 13 70 0.54
43 haberman 306 3 26 0.74
1524 vertebra-column 310 6 60 0.68
818 diggle_table_a2 310 8 79 0.53
1167 pc1_req 320 8 31 0.67
925 visualizing_galaxy 323 4 36 0.54
1011 ecoli 336 7 52 0.57
1048 jEdit_4.2_4.3 369 8 68 0.55
860 vinnie 380 2 13 0.51
1025 analcatdata_germangss 400 5 16 0.78
909 chscase_census2 400 7 70 0.51
1511 wholesale-customers 440 8 65 0.68
1498 sa-heart 462 9 79 0.65
814 chscase_vine2 468 2 18 0.55
724 analcatdata_vineyard 468 3 29 0.56
4329 thoracic_surgery 470 16 54 0.85
767 analcatdata_apnea1 475 3 20 0.87
884 fri_c0_500_5 500 5 50 0.50
750 pm10 500 7 70 0.51
886 no2 500 7 70 0.50
40690 threeOf9 512 9 9 0.54
335 monks-problems-3 554 6 15 0.52
333 monks-problems-1 556 6 15 0.50
949 arsenic-female-bladder 559 4 40 0.86
826 sensory 576 11 32 0.59
334 monks-problems-2 601 6 15 0.66
997 balance-scale 625 4 16 0.54
770 strikes 625 6 60 0.50
774 disclosure_x_bias 662 3 30 0.52
827 disclosure_x_noise 662 3 30 0.50
795 disclosure_x_tampered 662 3 30 0.51
931 disclosure_z 662 3 30 0.53
40981 Australian 690 14 80 0.56
1464 blood-transfusion-service 748 4 34 0.76
37 diabetes 768 8 73 0.65
1014 analcatdata_dmft 797 4 20 0.81
44268 anneal 898 38 104 0.54
50 tic-tac-toe 958 9 27 0.65
40693 xd6 973 9 9 0.67
799 fri_c0_1000_5 1000 5 50 0.50
45604 dummy 1000 6 60 0.73
43255 1StudentPerfromance 1000 7 43 0.52
741 rmftsa_sleepdata 1024 2 14 0.50
40702 solar-flare 1066 10 26 0.83
40706 parity5_plus_5 1124 10 10 0.50
934 socmob 1156 5 31 0.78
40680 mofn-3-7-10 1324 10 10 0.78
1462 banknote-authentication 1372 4 40 0.56
983 cmc 1473 9 35 0.57
40646 GAMETES_Epistasis_2-Way_2 1600 20 60 0.50
40649 GAMETES_Heterogeneity_20a 1600 20 58 0.50
991 car 1728 6 21 0.70
962 mfeat-morphological 2000 6 43 0.90
914 balloon 2001 1 10 0.76
772 quake 2178 3 30 0.56
40704 Titanic 2201 3 5 0.68
737 space_ga 3107 6 60 0.50
44127 phoneme 3172 5 49 0.50
3 kr-vs-kp 3196 36 38 0.52
871 pollen 3848 5 50 0.50
728 analcatdata_supreme 4052 7 24 0.76
720 abalone 4177 8 73 0.50
40983 wilt 4839 5 50 0.95
45039 compas-two-years 4966 11 34 0.50
44160 rl 4970 12 69 0.50
1460 banana 5300 2 20 0.55
803 delta_ailerons 7129 5 50 0.53
43922 mushroom 8124 22 109 0.52
807 kin8nm 8192 8 80 0.51
725 bank8FM 8192 8 73 0.60
816 puma8NH 8192 8 80 0.50
923 visualizing_soil 8641 4 31 0.55
819 delta_elevators 9517 6 47 0.50
44126 bank-marketing 10578 7 51 0.50
4534 PhishingWebsites 11055 30 46 0.56
45060 online_shoppers 12330 17 136 0.85
959 nursery 12960 8 26 0.67
1046 mozilla4 15545 5 40 0.67
44162 compass 16644 17 111 0.50
45558 Pulsar-Dataset-HTRU2 17898 8 80 0.91
45028 california 20634 8 80 0.50
823 houses 20640 8 80 0.57
43904 law-school-admission-bian 20800 10 53 0.68
843 house_8L 22784 8 75 0.70
45037 BitcoinHeist_Ransomware 24780 7 48 0.50
42493 airlines 26969 7 67 0.55
43900 amazon_employee_access 32769 9 90 0.94
44156 electricity 38474 8 68 0.50
44120 electricity 38474 7 61 0.50
137 BNG(tic-tac-toe) 39366 9 27 0.65
43901 click_prediction_small 39926 8 56 0.83
881 mv 40768 10 75 0.60
46554 Loan_Status 45000 13 94 0.78
45547 Cardiovascular-Disease-da 70000 11 49 0.50
45022 Diabetes130US 71090 7 41 0.50
40922 Run_or_walk_information 88588 6 60 0.50
Figure 24: Histogram of the number of instances and binarized features in the data sets considered in our experiments.
Additionally, we perform some experiments on data sets with more than 100,000 samples:
  • covertype (ID 44121) with 566,602 samples, 10 features, and 100 binarized features.
  • Higgs (ID 44129) with 940,160 samples, 24 features, and 240 binarized features.
Fig. 24 shows a histogram of the number of instances and the number of binarized features of the data sets considered in this paper.